### v1

# Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?

Abstract
Incorporating modular designs into neural networks demonstrates superior out-of-generalization, learning efficiency, etc. Existing modular neural networks are generally explicit because their modular architectures are pre-defined, and individual modules are expected to implement distinct functions. Conversely, recent works reveal that there exist implicit modular structures in standard pre-trained transformers, namely Emergent Modularity. They indicate that such modular structures exhibit during the early pre-training phase and are totally spontaneous. However, most transformers are still treated as monolithic models with their modular natures underutilized. Therefore, given the excellent properties of explicit modular architecture, we explore whether and how dense pre-trained transformers can benefit from emergent modular structures. To study this question, we construct Emergent Mixture-of-Experts (EMoE). Without introducing additional parameters, EMoE can be seen as the modular counterpart of the original model and can be effortlessly incorporated into downstream tuning. Extensive experiments (we tune 1785 models) on various downstream tasks (vision and language) and models (22M to1.5B) demonstrate that EMoE effectively boosts in-domain and out-of-domain generalization abilities. Further analysis and ablation study suggest that EMoE mitigates negative knowledge transfer and is robust to various configurations. Code is available at https://github.com/qiuzh20/EMoE
1 Introduction
Modularity attracts considerable attention from the artificial intelligence community (Auda & Kamel, 1999). Neural networks with modular designs, termed Modular Neural Networks (MNN), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), out-of-distribution (OOD) generalization abilities (Goyal & Bengio, 2020; Weiss et al., 2022), and data efficiency (Bengio et al., 2020). Typical MNNs are usually explicitly modular. Their modular structure is pre-defined and they are expected to achieve a divide-and-conquer solution for the given task. For example, Andreas et al. (2016) design separate visual and language functional modules and jointly train them to solve visual question-answering tasks. Among various MNNs, Mixture-of-Experts (MoEs) employ a conditional computation strategy where different submodules - so-called experts - are expected to be activated by different types of inputs. MoEs sees substantial success in various domains (Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and therefore becomes a widespread modular neural architecture.
Apart from developing explicit MNN,  Csordás et al. (2021); Agarwala et al. (2021) study whether standard neural networks become implicitly modular after training and discover spontaneously emerged modular structure in small-scale CNNs and LSTMs. For more complicated large-scale pre-trained transformers, initial observations (Zhang et al., 2022b; Li et al., 2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs). Specifically, they find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process. However, sparsity does not imply modularity. Sparsely activated FFNs neurons do not ensure to have remarkable functional division. Therefore,  Zhang et al. (2023) further utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs. They observe a strong correlation between activation and specific tasks, further discovering clear function-based neuron grouping of the model and summarizing such phenomenon as Emergent Modularity.
Figure 1: (a) Existing literatures (Geva et al., 2021; 2022) suggest that the FFNs in transformers can be viewed as key-value memories. They regarded the input as a query, the first layer as keys, and the second as values. Given an input, keys are sparsely activated (marked in red). Most of the values don’t impact the output. (2) The FFNs block can be partitioned into experts by clustering keys. (3) Afterward, experts’ keys averages are used as the gating weights. The inner product between x and gating weights are used to select experts.
However, contemporary transformers are still employed as monolithic models, with their emergent modularity underutilized. Motivated by the advantages of explicit modularity, we explore whether and how standard pre-trained transformers could benefit from emergent modular structures. Technically, we externalize the emergent modularity into the explicit MoEs, i.e., Emergent Mixcture-of-Experts (EMoE). As shown in Figure 1, we first (1) utilize a clustering-based method to decompose the FFNs into experts and then (2) use the Average-Key-Gating ($\operatorname{avg-k}$ gating) to route the input to corresponding experts. Both two steps do not require extra training processes and data.
Then we fine-tune EMoE on a wide range of downstream tasks to investigate if standard pre-trained transformers can benefit from its ermgent modular structure. We evaluate EMoE following different configurations: (1) tasks from various modalities: vision (Domainbed (Gulrajani & Lopez-Paz, 2021)) and language (GLUE (Wang et al., 2019) and GLUE-X (Yang et al., 2023)); (2) different pre-trained dense transformers: ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), and GPT2 (Radford et al., 2019), ranging from 22M to 1.5B parameters. (3) different evaluation settings: in-domain (ID) evaluation and OOD evaluation; (4) different fine-tuning methods: full fine-tuning and parameter-efficient tuning (Hu et al., 2022). In total, we fine-tune 1785 models.
Refer to caption Refer to caption
Figure 2: EMoE achieves stronger ID and OOD performances compared with baselines in both full-fine-tuning (Left, average accuracy) and LoRA tuning settings (Hu et al., 2022) (Right, average accuracy for ID and Friedman rank (Friedman, 1940) for OOD). Results are summarized from Section 4.
Our results are summarized in Figure 2. We find that fine-tuning EMoE achieves significant improvement compared with the vanilla fine-tuning and competitive performance with strong baseline GMoE (Li et al., 2023), demonstrating that dense pre-trained transformer can indeed benefit from emergent modular structures to achieve better generalization performance. We further find that EMoE can deactivate neurons with negative transfer effects, thus improving downstream performance. Meanwhile, our ablation studies indicate that EMoE’s insensitivity to various hyper-parameter configurations and higher data efficiency of up to 20%.
2 Related Work
Modular neural networks has various advantages, including interpretability (Pfeiffer et al., 2023), scalability (Chowdhery et al., 2022), multi-task learning abilities (Chen et al., 2023a), and OOD generalization abilities (Goyal et al., 2021; Li et al., 2023). MoEs (Szymanski & Lemmon, 1993) is currently regarded as a standard framework for developing modular neural networks (Shen et al., 2023b). Most MoEs models are developed by training explicit MoEs architecture from random intialization (Shazeer et al., 2017; Fedus et al., 2022b; Zhang et al., 2022a), which usually causes load imbalance issues (Zoph et al., 2022; Fedus et al., 2022a; Chi et al., 2022). Some recent works employ already-trained dense models as a good initialization (Nie et al., 2021). For example, GMoE (Li et al., 2023) and Upcycling Komatsuzaki et al. (2023) copy the FFNs from a trained dense network to continue pre-training or transfer learning. EMoE is also based on pre-trained dense models. However, unlike existing methods that often significantly increase in parameters (GMoE increases 50% parameters) or tend to be less effective (Upcycling extends T-5 Base (248M) to 2B parameters but still underperforms the T5-Large (783M)), EMoE achieves significant improvements without introducing extra training processes or parameters. This is primarily attributed to the heterogeneity of its experts and gating mechanisms, as we demonstrate in Section  5 and 6.
In another research line,  Hod et al. (2021); Csordás et al. (2021) explore modular structures in CNNs and LSTMs. Some works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated dense pre-trained transformers. Based on their observations,  Zhang et al. (2023) recently explore modularity in pre-trained transformer FFNs. In this work, we adopt clustering-based expert construction and $\operatorname{avg-k}$ gating from MoEfication (Zhang et al., 2022b), while our motivation differs. Seeking to preserve the original functionality with fewer activated neurons, they split every layer in the dense model into MoEs. However, the results are worse than the dense model by more than ten points after splitting (Table 2 in their original paper). This is because the clustering-based decomposition and $\operatorname{avg-k}$ gating cannot precisely select the activated neurons. Without any tuning, these errors accumulate with MoEs layers and lead to severe performance dropping. Therefore, Zhang et al. (2022b) mainly focus on more complicated learning-based construction methods. Differently, we treat these decomposed units as heterogeneous modules to exploit their modular properties in the framework of MoEs, and observe significant improvements over monolithic models.
3 Methodology
3.1 Preliminaries
Transformer Feed-Forward Networks The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{d\times h}$, $\mathbf{V}\in\mathbb{R}^{h\times d}$, and a non-linear activation function $\sigma(\cdot)$. For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as follows:
$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}.$ (1)
More precisely, for each row $\mathbf{K}_{i,:}$ and column $\mathbf{V_{:,i}}$, the Equation 1 can be rewritten as:
$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K_{i,:}})\cdot\mathbf{V_{:,i}}$ (2)
Regarding each row in $\mathbf{K}$ as a key vector and each column in $\mathbf{V}$ as the value vector, the output of an FFNs network can be viewed as a weighted sum of values based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K}^{T})$.
Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module. Supposing there are $N$ experts: $\{\operatorname{FFN}_{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$. Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function. Given an input $\mathbf{x}\in\mathbb{R}^{h}$, let $\mathbf{y}_{n}=\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ be the output of the $n$-th expert, and then the output $\mathbf{y}\in\mathbb{R}^{h}$ of can also be summarized as the weighted sum of the output from all experts:
$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)y_{n},$ (3)
When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation.
3.2 Emergent Mixture-Of-Expert
The ideal EMoE should reflect the implicit modular nature of dense pre-trained transformers using the form of MoEs. A preferred approach to constructing EMoE should avoid introducing additional parameters, training, and data, which may result in undesired biases and are impractical. Although more sophisticated constructing methods can be proposed, we adopt a method from MoEfication Zhang et al. (2022b). The method contains two steps: (1) clustering-based expert construction and (2) $\operatorname{avg-k}$ gating. It is simple and fulfills our needs, as demonstrated in our experiments.
Clustering-based Experts Construction. Since key-value pairs with similar key vectors tend to be co-activated, we split them into separate experts according to their key vectors. Specifically, given a trained FFNs layer $\operatorname{FFNs}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen & Fränti, 2014) to partition all key vectors $\mathbf{K}$ into $N$ experts on average, so each group contains $\frac{d}{N}$ key-value pairs (usually $d=4h$). Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract key-value pair $(\mathbf{K_{j,:}},\mathbf{V_{:,j}})$ to form the $i$th expert $\operatorname{FFN}(\cdot;\mathbf{K}_{i},\mathbf{V}_{i})$, as depicted in Figure 1(b). After that, the computation of each expert proceeds as Equation 1.
Avg-k Gating. Given an input $\mathbf{x}$ and $N$ experts, a qualified gate for EMoE should route the input $\mathbf{x}$ to the experts who contribute most to the model’s output. We construct the gating module by averaging each expert’s keys. The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $n$th column $\mathbf{G}_{,:n}$ can be calculated as follows:
[…]
And then the corresponding gating score for $i$-th expert is:
$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G}^{T};k,\text{dim=1})\\ &0~{}~{}\text{ else}\end{cases}$ (5)
where $\operatorname{Top-K}(\cdot;k,\text{dim})$ return indices of $k$ largest element of the given input along a given dimension. Using 0,1 score, $\operatorname{avg-k}$ gating reduces the weights’ influence in Equation 3). Notably, observe:
$\mathbf{x}\cdot\mathbf{G}_{:,n}^{T}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{n},\text{dim}=0)=\mathbf{x}\cdot\frac{N}{h}\sum_{j\in E_{i}}\mathbf{K}^{n}_{j,:}=\frac{N}{h}\sum_{j}a_{j}.$ (6)
A larger value of gating score $g_{i}$ implies more activated keys within the corresponding expert. Consequently, the expert could potentially contribute more to the output $\mathbf{y}$ for input $\mathbf{x}$. During downstream tuning, gating weights are updated along with the FFNs parameters using Equation 4.
3.3 Discussion
Though adopting a method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. However, the method introduces deviations from FFNs as it can mistakenly eliminate useful neurons. As these deviations accumulated across layers, the performances severely drop by over 10 points on original tasks. Therefore, they focus on more precise expert construction and gating to preserve performances. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. We summarize their main differences in Appendix A.1. As one of our most related work, GMoE (Li et al., 2023) introduce MoEs by copying the trained FFNs. They further demonstrate introducing MoEs every 2 layers leads to significant degradation. Accordingly, we only introduce EMoEs in certain layers since efficiency isn’t our primary concern. We also conduct ablation studies about MoEfying every 2 layers and get consistent results, more details are in Appendix A.4.2 Table 14.
4 Experiments
[…]
Baselines On Domainbed, our baselines include (1) vanilla ViT, as it is a strong OOD baseline under fair configurations suggested by Gulrajani & Lopez-Paz (2021). (2) GMoE (Li et al., 2023), which is the latest state-of-the-art. GMoE constructs MoEs by replicating FFNs from the two-to-last and fourth-to-last transformer blocks of a pre-trained ViT and therefore is the most relevant to EMoE. For language tasks, besides vanilla backbone and GMoE, we implement (3) noisy tuning (Wu et al., 2022), which also improves adaptation for free by adding uniform distribution noise proportional to the standard deviation of the pre-trained weights before tuning. (4) EMoE-learn, an ablation method, where the gating function is learned (same as GMoE) during fine-tuning instead of employing $\operatorname{avg-k}$ gating. This helps us better understand the effect of $\operatorname{avg-k}$ gating.
Table 1: Overall OOD performances with 3 selection criteria. All the reported results are obtained following the Domainbed codebase. The best result is highlighted in bold. In cases where results are the same, the best result is determined by the smallest standard deviation. EMoE demonstrates comparable results to GMoE.
Results with ViT-small (22M) backbone
Results with ViT-base (86M) backbone
Algorithm PACS VLCS
Office
Terra
Avg
Train-validation selection criterion
ViT 86.9 79.7 73.0 44.0 70.90
GMoE 87.7 79.6 73.1 45.4 71.45
EMoE-learn
87.2 79.6 72.5 46.1 71.35
EMoE 87.8 79.5 73.1 45.9 71.58
Leave-one-domain-out selection criterion
ViT 86.1 79.7 73.3 45.0 71.03
GMoE 86.5 80.5 73.1 45.3 71.35
EMoE-learn
86.8 79.6 72.6 45.8 71.20
EMoE 86.8 80.6 73.3 46.1 71.70
Test-domain selection criterion
ViT 86.5 78.2 73.1 44.0 70.45
GMoE 87.2 79.0 73.4 45.3 71.23
EMoE-learn
87.4 79.1 72.8 45.4 71.18
EMoE 87.6 79.0 73.3 45.5 71.35
Algorithm PACS VLCS
Office
Terra
Avg
Train-validation selection criterion
ViT 89.1 80.7 77.2 47.3 73.58
GMoE 90.0 80.4 77.0 49.2 74.15
EMoE-learn
89.8 80.6 76.5 48.7 73.90
EMoE 89.4 80.7 77.3 48.5 73.98
Leave-one-domain-out selection criterion
ViT 88.9 80.8 77.5 46.1 73.33
GMoE 89.3 81.0 76.7 50.1 74.28
EMoE-learn
89.3 81.2 76.5 50.5 74.38
EMoE 89.6 81.6 77.4 50.0 74.65
Test-domain selection criterion
ViT 88.8 79.0 77.2 46.7 72.93
GMoE 89.7 79.0 77.0 48.8 73.63
EMoE-learn
89.7 79.7 76.6 48.7 73.68
EMoE 89.7 79.7 77.5 48.8 73.93
Results with ViT-base (86M) backbone
4.1 Full Fine-tuning Performance
We test the EMoE’s OOD performance on Domainbed. Domainbed provides comprehensive vision OOD evaluations (one result is aggregated with 30 experiments), and the outcomes vary marginally. Moreover,  Gulrajani & Lopez-Paz (2021) indicates that vanilla full fine-tuning with fair hyper-parameter search is a strong baseline compared with specifically designed methods like Invariant Risk Minimization (Arjovsky et al., 2019). More dataset details are in appendix A.2.1. According to Table 1: (1) Overall, EMoE outperforms ViT and GMoE (except ViT-base Train-validation setting, upper right). (2) Compared with EMoE, EMoE-learn incorporates a learned gate. While it surpasses $\operatorname{avg-k}$ gating in certain scenarios (ViT-small Terra Train-validation), it can also lead to a performance drop compared with vanilla ViT. Its overall performance is lower than EMoE. (3) In tasks where the dense model is strong (like Office), EMoE performs better than other MoEs methods. One possible reason is that the $\operatorname{avg-k}$ gating reduces the influence of gating weights ($g_{n}(\mathbf{x};\mathbf{G};k)$ in Equation 3), making it more like the dense model in such scenarios.
Table 2: ID performance on GLUE tasks with different backbones and algorithms. All the reported results are obtained from 3 independent experiments. The average accuracy (Avg) is reported along with the relative improvement compared to the baseline. The best result is highlighted in bold.
Backbone
Algorithm
MRPC CoLA RTE STSB SST2 Avg
BERT Base baseline 88.45 60.67 68.95 87.87 91.97 79.582
noisy tuning
88.43 61.79 71.36 88.27 92.32 80.43(+0.85)
GMoE 88.63 61.25 70.28 88.63 92.28 80.21(+0.63)
EMoE-learn
89.05 62.46 70.40 88.47 92.58 80.59(+1.01)
EMoE 89.45 61.55 69.68 88.71 92.89 80.46(+0.87)
BERT Large baseline 89.82 65.41 74.89 89.87 93.50 82.70
noisy tuning
90.42 64.75 73.41 90.05 93.65 82.46(-0.24)
GMoE 91.24 64.90 74.24 90.00 93.58 82.79(+0.09)
EMoE-learn
90.57 65.51 74.72 90.22 93.73 82.95(+0.25)
EMoE 90.74 65.79 76.17 90.31 93.58 83.32(+0.62)
GPT2 Small baseline 84.46 47.07 67.15 86.29 92.13 75.42
noisy tuning
84.15 46.16 67.51 86.09 92.13 75.21(-0.21)
GMoE 85.07 47.77 67.51 86.57 92.35 75.85(+0.43)
EMoE-learn
85.73 47.24 67.99 86.66 92.35 75.99(+0.57)
EMoE 85.40 48.00 68.95 86.64 92.70 76.34(+0.92)
We evaluate EMoE’s ID performance on 5 GLUE tasks. According to Table 2: (1) On average, EMoE and EMoE-learn outperform other baselines. (2) Among the two methods that do not introduce additional parameters, EMoE significantly outperforms noisy tuning. (3) EMoE provides stable improvements over baselines across different settings, demonstrating its generality.
4.2 LoRA-Tuning performance
Table 3: ID and OOD performance of EMoE and baseline models. All the reported results are obtained from 3 independent experiments. OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization. The best result is highlighted in bold.
Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
OOD
BERT-Large (340M Parameters)
LoRA 89.97 63.40 72.92 90.51 93.16 92.20 87.21 85.40 84.35 4.86
Block 89.34 62.10 71.96 90.39 93.35 92.04 88.45 86.20 84.23(-0.12) 4.95
GMoE 89.45 63.80 72.56 90.29 93.85 92.32 87.99 85.92 84.52(+0.18) 4.04
EMoE-block
89.77 63.25 71.60 90.31 93.69 92.09 88.08 86.21 84.38(+0.03) 5.89
EMoE-learn
89.87 64.00 71.36 90.48 93.65 92.40 87.55 85.62 84.37(+0.02) 4.66
EMoE 90.85 65.33 75.21 90.43 93.50 92.23 87.74 85.43 85.09(+0.74) 4.37
GPT2-XL (1.5B Parameters)
LoRA 86.83 60.88 78.70 89.07 95.18 91.84 87.41 86.93 84.61 5.61
Block 86.59 61.18 79.78 89.08 95.45 91.88 87.71 86.95 84.83(+0.22) 5.13
GMoE 87.02 62.81 79.78 89.21 95.41 92.18 89.10 87.17 85.34(+0.73) 4.33
EMoE-block
87.86 62.88 80.05 89.18 95.49 92.10 89.69 86.87 85.52(+0.91) 5.71
EMoE-learn
87.93 61.50 79.90 89.48 95.18 92.33 89.71 87.00 85.38(+0.77) 4.40
EMoE 87.75 62.27 80.02 89.37 95.41 92.10 89.58 87.06 85.45(+0.84) 3.88
With the increasing scale of pre-trained models, parameter-efficient tuning (Houlsby et al., 2019) gets popular. Observing weaker results with LoRA-tuned ViT-Large on Domainbed compared to ViT-Base, we omit LoRA tuning on Domainbed. GLUE-X (Yang et al., 2023) is a recently introduced OOD dataset for the GLUE benchmark. Therefore, we assess EMoE’s ID performance on GLUE and OOD performance on GLUE-X with the standard LoRA tuning. Notably, LoRA weights are only added in attention, with all other weights frozen. Therefore, the only difference between EMoE and its dense counterpart is FFNs activation. For the OOD metric, we follow GLUE-X (Yang et al., 2023) and employ the Friedman rank (Friedman, 1940) $\text{rank}_{f}=\frac{1}{n}\sum_{i=1}^{n}\text{rank}_{i}$. For each method under the same backbone, $\text{rank}_{i}$ is produced based on the best result and the average result for each dataset. For the 13 OOD tasks used in GLUE-X, each method generates 26 $\text{rank}_{i}$ values. The OOD results presented in Table 3 represent the mean of all these $\text{rank}_{i}$ values. The original results of each task can be found in the Appendix A.4.1. For algorithm configurations, since GMoE copies the FFNs of the pre-trained model to form the MoEs, it is meaningless if the MoEs parameters aren’t tuned. Thus, in Table 3, we conduct experiments with LoRA tuning for GMoE and also tune the transformer block where the original FFNs is replaced. For comparison, we also (1) LoRA-tune dense model and fine-tune the transformer block at the corresponding layer (denoted “Block”); (2) LoRA-tune EMoE with the transformed block also updated (denoted “EMoE-block”).
According to Table 3, EMoE continues to enhance the downstream performance in LoRA tuning: (1) EMoE demonstrates significant enhancements compared to LoRA and doesn’t introduce additional trainable parameters or procedures. Notably, EMoE achieves comparable results with GMoE on BERT-large and outperforms it on GPT2-XL. (2) When the blocks are tuned, EMoE also brings improvements (EMoE-block vs. Block) (3) Consistent with the full-finetuning findings, EMoE exhibits higher stability than EMoE-learn and delivers superior overall results. (4) Besides vision, MoEs also improve OOD performance in language tasks (GMoE vs. Block, EMoE vs. LoRA).
5 Analysis
5.1 How does EMoE improve fine-tuning performance?
Refer to caption     Refer to caption
Figure 3: Left: ID and OOD results compared with LoRA for validating EMoE’s training & inference effects. Right: sparse activated training results with different expert selections.
EMoE achieves notable improvements in both ID and OOD scenarios. However, it is non-trivial that simply transforming the pre-trained FFNs into MoEs before fine-tuning can yield such benefits (especially in LoRA tuning where the MoEs part is totally not updated). Therefore, we investigate the reasons behind the enhanced fine-tuning performance. To decouple the impact of updating expert parameters, we focus on LoRA tuning. We use BERT-Large as the backbone model.
EMoE benefits LoRA weights learning instead of influencing inference. EMoE and dense model only differ in FFNs activations; such difference might (1) directly impact outputs during testing and (2) influence the parameter updating during training. In light of this, we propose two variants and compare them with vanilla LoRA tuning and LoRA tuning with EMoE: (a) LoRA2EMoE: splitting a LoRA-tuned dense model into its MoEs counterpart. If the test results differ from the original ones, we can infer that the sparse activation has an impact during testing. (b) EMoE2LoRA: merging the LoRA weights tuned with EMoE back into the dense. If no changes occur, it implies that the primary factor is the parameter updating. According to Figure 3 Left, doing sparse activation during testing has little impact (LoRA2EMoE). But when merging LoRA weights tuned with EMoE back into the dense, the performance remains significantly better than the vanilla LoRA tuning (EMoE2LoRA vs. LoRA) and is almost identical to EMoE. For full results, please refer to Appendix A.4 Table 11.
EMoE masks neurons with negative transfer impacts. The only difference between EMoE and vanilla LoRA tuning is EMoE blocks some activated neurons during training by $\operatorname{Top-k}$ expert selection. Based on these, we hypothesize that EMoE’s effects stem from preventing negative knowledge transfer from blocked neurons. Therefore, we investigate whether there are such negative transfers. Specifically, we study the following expert selection variants: (1) Bottom-k: select $k=16$ experts who get the lowest scores; (2) Not-top-k: select experts who are not among the top-k experts. These variants are evaluated across 6 tasks from GLUE. The averaged outcomes are in Figure 3 Right. Full results can be found in Appendix A.4 Table 12. LoRA tuning results with “Bottom-k” and “Not-top-k” expert selections are worse than vanilla LoRA tuning. One may attribute this drop to the reduced number of employed neurons. To further assess this, we compare “Top-k” with “Bottom-k”, where an equal number of neurons are used. Notably, “Bottom-k” significantly lags behind “Top-k”. This further corroborates that masked neurons have negative transfer effects.
5.2 Ablation studies
Figure 4: Left: Activations of neurons in FFNs a pre-trained ViT (Dosovitskiy et al., 2021). Right: By clustering the keys in the FFNs and rearranging the action values accordingly, activation patterns of neurons emerge. These co-activated patterns correspond to the modular structure within the pre-trained dense model. With appropriate partitioning, we can transform dense networks into modular networks.
Table 4: Clustering-based and random expert constructing results
Top-k Bottom-K
Cluster +0.92 -0.48
Random -0.11 -0.34
Expert Constructing Methods Figure 4 demonstrates that clustering can decompose modular components within the dense model. To provide further evidence that the EMoE’s improvements stem from leveraging modular features rather than just sparse activation, we compare the results of (1) clustering-based expert construction and (2) random construction within the same setting of Section 3. The relative changes in averaged outcomes compared to the dense baseline are shown in Table 4, while full results can be found in the appendix A.4, Table 12. It’s noteworthy that while cluster top-k exhibits a significant improvement over dense, random top-k is conversely worse than dense baseline. This suggests that random construction can negatively impact gating. Moreover, considering selecting weights with negative transfer under bottom-k selections, it’s observed that cluster bottom-k also achieves lower results. In summary, clustering-based methods can externalize implicit modularity within the dense model. Within suitable frameworks like MoEs, such modularity can facilitate downstream tuning.
Figure 5: Results of 3 tasks for different experts splittings and expert select fractions (top-k / number of experts). ‘e=16’ and ‘e=32’ mean splitting FFNs into 16 and 32 experts, respectively.
Expert Constructing Configurations Beyond the settings detailed in the main results section, which are based on $N=64$ experts and top-k $\in\{16,32,48\}$, we also present specific scenarios where $N\in\{16,32\}$, and top-k varies within $\{2,4,8,16\}$ in Figure 5. Notably, in each of these settings, (1) EMoE consistently surpasses the dense model, illustrating its insensitivity to hyper-parameters. (2) On average, $\operatorname{avg-k}$ gating exhibits superior performance than learned gating. Though learned gating (EMoE-learn) outperforms $\operatorname{avg-k}$ gating in a few specific settings (Figure 5 (b) and (e)). This is consistent with the earlier results in Section 4. Regarding how many EMoE layers should be introduced, our findings align with those discovered in GMoE, indicating that only a limited number of layers can be converted into the EMoE layer. If excessive EMoE layers are introduced, performance deteriorates. Taking GPT2-XL (48 layers) as an example, when introducing EMoE every 2 layers in the latter half, the performance averaged across 5 GLUE tasks (79.36) matches that of the dense model (78.87). But when adopting EMoR every 2-layer for the entire model, the performance lags slightly behind that of the dense model (78.17) but surpasses EMoE-learn (75.87). For additional configurations, please refer to the appendix A.4.2 Table 14.
Figure 6: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs. learned) and expert types (modules from dense vs. repetitions of dense). The vertical axis illustrates training steps (top-down arrangement signifies begin-end); the horizontal axis represents expert selection frequency within 1K steps (deeper color implies a higher frequency). (a), (b) and (c) correspond to EMoE, EMoE-learn, and GMoE.
Expert Selection Changes During Training To further understand $\operatorname{avg-k}$ gating and learned gating, we visualize expert selections of GPT2-XL during fine-tuning on 6 tasks with 16 experts. In Figure 6, we showcase the results for the largest dataset QNLI among them. Full results are available in the appendix A.5. Our observations are: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating (Figure 6 a vs. b). This might mitigate data inefficiency from gating inconsistencies across different stages of training (Zuo et al., 2022). (3) EMoE, with its heterogeneous experts, exhibits better load balancing than GMoE (Figure 6 b vs. c). In GMoE, all experts share identical initialization, whereas in EMoE, the experts are derived from FFNs with implicit modularity. This also suggests a good initialization can facilitate MoEs learning similar to EvoMoE (Nie et al., 2021).
Figure 7: Average performance of EMoE with different proportions of training data.
Performance Across Different Training set Volumes. Previous research has indicated that modular architectures offer improved data efficiency (Bengio et al., 2020). Therefore, we conducted experiments with GPT2-XL on six tasks using varying proportions of original training data, and the results for all tasks are presented in Figure 7. It can be observed that EMoE consistently outperforms the dense across different data factions. EMoE achieves superior results even when using less than 20% of the data. On SST2, only using 50% data, EMoE shows comparable performance to the dense. More details can be found in the Appendix A.4.2 Table 15. This further underscores the benefits of incorporating modular structures.
6 Conclusion, Limitations and Future Works
Conclusion: In this work, we validate that exploiting the emergent modular structures in dense pre-trained transformers improves downstream task ID and OOD performances. One possible reason is the modular structure can alleviate negative transfer effects presented in the pre-trained model. We hope our findings could deepen the understanding of neural networks’ modularity, further helping the community develop more sophisticated modular neural architectures and utilizing existing pre-trained models. Limitations: We would like to point out that EMoE is not designed to outperform the SOTA tuning methods and OOD generalization algorithms. Instead, the goal is to demonstrate that dense pre-trained transformers can benefit from emergent modular structures. Our research findings have not been validated on more challenging tasks (e.g., Mathematical Reasoning (Imani et al., 2023)) or foundation models like LLaMA (Touvron et al., 2023) due to computational constraints. While our evaluation was primarily conducted on models with a maximum parameter count of 1.5B, it is evident that across the spectrum of parameter sizes, ranging from 22M to 1.5B, EMoE consistently yields improvements, showing the potential to scale up. Future works: Subsequent research could delve into better methods for modular decomposition and investigate the modularity in foundation models. Benefiting from EMoE’s ability to avoid negative transfer, future researchers can also exploit the emergent modularity in settings such as multi-task learning and continual learning. Moreover, it is observed that modularity structures emergent and remain stable after approximately one-fourth of the pre-training phase (Li et al., 2022; Zhang et al., 2023). Future investigations could investigate leveraging the modularity of dense models during the pre-training process.
7 Reproducibility
In the Experiment Configuration Section 4, we first introduce the additional configurations introduced by EMoE relative to the Dense Model, as well as the baseline settings. Subsequently, we provide high-level experiment settings for each experimental group (Domainbed at 4.1, full fine-tuning GLUE tasks at section 2, LoRA tuning at  3). In Appendix A.2, we provide detailed information on the datasets used, the codebase, and the evaluation metrics. We also outline more general configurations, such as learning rates and batch sizes for the respective tasks. Beyond the main paper, we include the original code for the experiments and log files for certain experimental results in the supplementary materials. Each section of code includes a README.md file that explains the experimental settings required to replicate the results.
Appendix A Appendix
A.1 Compare with Moefication
Table 5: Comparison between EMoE and Moefication
Aspect EMoE Moefication
Motivation
Modularity
Sparsity
Goal
Transfer learning effective
Inference efficiency
MoEs Layers
Typically 1 or 2 layers
Applies changes across all layers
Scenarios
Down-stream tasks
Original task
Results
Significant improvement without adding parameters
Improved inference efficiency but performance drop
Though adopting the MoEs construction method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. Accordingly, We only introduce EMoEs in certain layers according to GMoE, since efficiency is not our primary concern. These differences are summarized in Table 5.
A.2 Datasets and Evaluation Metrics
A.2.1 Domainbed
Table 6: Used dataset information from Domainbed
Dataset PACS VLCS OfficeHome TerraInc
#Domains 4 4 4 4
Classes 7 5 65 10
Images 9,991 10,729 15,588 24,788
The four datasets (PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), and Terra Incognita (Beery et al., 2018)) are selected from Domainbed. Each dataset comprises 4 distinct domains(PACS: {art, cartoons, photos, sketches}, VLCS: {Caltech101, LabelMe, SUN09, VOC2007}, Office-Home: {art,clipart, product, real}, and Terra Incognita: {L100, L38, L48, L46}). Within a single training, one or two domains’ data are sequentially designated for OOD evaluation. For example, when training on PACS, {art, cartoons} could be selected as ID training data, while {photos, sketches} are designated for OOD testing. This configuration results in a total of $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset. Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited accesses to test domains and selecting based on these results. The final results are aggregated with selection criteria provided by Domainbed [^1] https://github.com/facebookresearch/DomainBed. As a result, even a variation of 0.1 in the benchmark outcomes signifies a significant improvement.
In our experiments, all the hyper-parameters, like training steps, learning rates, and weight decay, except those related to MoEs, strictly follow GMoE.
A.2.2 GLUE
Each task involves one to four OOD tasks from GLUE-X (Yang et al., 2023), resulting in 13 OOD tasks in total. To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani & Thummar, 2023) from GLUE-X for OOD testing.
In the full fine-tuning, to ensure convergence and reduce randomness, we train all models 10 epochs across 3 random seeds on each task. Each experiment does a hyper-parameter search on learning rates on [2e-5, 3e-5, 5e-5] as suggested by BERT (Devlin et al., 2019). The training batch size is 32. In the LoRA tuning, following LoRA (Hu et al., 2022) that tunes models with more epochs and larger learning rates than standard full fine-tuning, all models are trained 20 epochs on small and medium datasets and 5 epochs on large ones (like QNLI, MNLI, QQP). The learning rate is searched in [2e-4, 3e-4, 5e-4]. All methods are implemented with LoRA_rank=8 and LoRA_alpha=16. The training batch size is 16 due to a larger model size. Other settings like max_lengt following the codebase from hugging face [^2] https://github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification. After training on GLUE, we directly test the selected models on GLUE-X with the data from the official repo [^3] https://github.com/YangLinyi/GLUE-X.
Table 7: Language tasks and corresponding ID and OOD datasets.
Task ID-dataset size OOD-dataset size
Paraphrase MRPC 4,076 Twitter 16,777
QQP 404,301 Twitter 16,777
MRPC 4,076
Linguistic Acceptability CoLA 9,594 Grammar Test 304,277
Textual Entailment RTE 2,768 SciTail 26,527
HANs 60,000
Textual Similarity STSB 7,128 SICK 9,840
Sentiment Analysis SST2 68,223 IMDB 50,000
Yelp 598,000
Amazon 4,000,000
Flipkart 205,041
Question Answering NLI QNLI 110,206 NewsQA 119,525
Natural Language Inference MNLI 412,313 SICK 9,840
A.3 Computation Cost and Memory Usage
Theoretically, EMoE does not introduce additional parameters compared to its dense counterpart. Although it adds computation in the gating portion within the MoEs layer, it omits a substantial amount of computation within the FFNs layer. For instance, the computation in the gating portion is of the order of $h\times N$, where h represents the model’s hidden size, and N is the number of experts. In contrast, the complete computation in the FFNs layer is of the order of $(h\times h\times 4h)\times 2$, and sparse activations can reduce more than a quarter of this computation. Since $N\ll h$, theoretically, using EMoE within a single block should accelerate the forward pass of the model. However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the dense model. Additionally, the memory usage is also slightly higher, by less than 5%, compared to the dense model (16295MB v.s. 15865MB on CoLA, LoRA tuning GPT2-XL).
Each individual experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
A.4 Tabular Results
A.4.1 Full Tables in Full Fine-tuning with Standard Deviation
In this section, we present the mean and variance of experiments conducted with three different random seeds. The Domainbed results are demonstrated in Table 8, full fine-tuning results are in Table 9, LoRA tuning results are in Table 10.
Table 8: Overall out-of-domain performances with different selection criteria. All the reported results are obtained from three independent experiments conducted following the Domainbed codebase. The best result is highlighted in bold.In cases where results are the same, the best result is determined by the smallest standard deviation. EMoE demonstrates comparable results to GMoE.
Results with ViT-small backbone
Results with ViT-base backbone
Algorithm PACS VLCS OfficeHome TerraInc Avg
train-validation selection criterion
ViT 86.9
$\pm$0.2
79.7 $\pm$0.4 73.0
$\pm$0.2
44.0
$\pm$1.1
70.90
GMoE 87.7
$\pm$0.2
79.6
$\pm$0.4
73.1
$\pm$0.3
45.4
$\pm$0.3
71.45
EMoE-learn
87.2
$\pm$0.4
79.6
$\pm$0.2
72.5
$\pm$0.2
46.1 $\pm$0.4 71.35
EMoE 87.8 $\pm$0.2 79.5
$\pm$0.4
73.1 $\pm$0.2 45.9
$\pm$0.3
71.58
leave-one-domain-out selection criterion
ViT 86.1
$\pm$0.6
79.7
$\pm$0.4
73.3 $\pm$0.1 45.0
$\pm$0.5
71.03
GMoE 86.5
$\pm$0.3
80.5
$\pm$0.2
73.1
$\pm$0.3
45.3
$\pm$0.6
71.35
EMoE-learn
86.8 $\pm$0.0 79.6
$\pm$0.3
72.6
$\pm$0.2
45.8
$\pm$0.6
71.20
EMoE 86.8
$\pm$0.1
80.6 $\pm$0.4 73.3
$\pm$0.2
46.1 $\pm$0.6 71.70
test-domain selection criterion
ViT 86.5
$\pm$0.4
78.2
$\pm$0.4
73.1 $\pm$0.2 44.0
$\pm$0.5
70.45
GMoE 87.2
$\pm$0.4
79.0
$\pm$0.3
73.4 $\pm$0.2 45.3
$\pm$0.4
71.23
EMoE-learn
87.4
$\pm$0.2
79.1 $\pm$0.3 72.8
$\pm$0.1
45.4
$\pm$0.6
71.18
EMoE 87.6 $\pm$0.5 79.0
$\pm$0.2
73.3
$\pm$0.0
45.5 $\pm$0.1 71.35
Algorithm PACS VLCS OfficeHome TerraInc Avg
train-validation selection criterion
ViT 89.1
$\pm$0.0
80.7 $\pm$0.1 77.2
$\pm$0.1
47.3
$\pm$0.8
73.58
GMoE 90.0 $\pm$0.3 80.4
$\pm$0.6
77.0
$\pm$0.1
49.2 $\pm$1.1 74.15
EMoE-learn
89.8
$\pm$0.2
80.6
$\pm$0.2
76.5
$\pm$0.1
48.7
$\pm$0.5
73.90
EMoE 89.4
$\pm$0.4
80.7
$\pm$0.2
77.3 $\pm$0.1 48.5
$\pm$0.5
73.98
leave-one-domain-out selection criterion
ViT 88.9
$\pm$0.4
80.8
$\pm$0.3
77.5 $\pm$0.1 46.1
$\pm$0.6
73.33
GMoE 89.3
$\pm$0.6
81.0
$\pm$0.3
76.7
$\pm$0.1
50.1
$\pm$1.1
74.28
EMoE-learn
89.3
$\pm$0.2
81.2
$\pm$0.1
76.5
$\pm$0.1
50.5 $\pm$0.2 74.38
EMoE 89.6 $\pm$0.2 81.6 $\pm$0.2 77.4
$\pm$0.1
50.0
$\pm$1.1
74.65
test-domain selection criterion
ViT 88.8
$\pm$0.7
79.0
$\pm$0.5
77.2
$\pm$0.0
46.7
$\pm$0.4
72.93
GMoE 89.7
$\pm$0.5
79.0
$\pm$0.3
77.0
$\pm$0.1
48.8 $\pm$0.4 73.63
EMoE-learn
89.7 $\pm$0.4 79.7 $\pm$0.2 76.6
$\pm$0.1
48.7
$\pm$0.3
73.68
EMoE 89.7 $\pm$0.4 79.7 $\pm$0.2 77.5 $\pm$0.1 48.8
$\pm$0.6
73.93
Results with ViT-base backbone
Table 9: Results on GLUE tasks with different backbones and algorithms. All the reported results are obtained from 3 independent experiments. The average accuracy (avg) is reported along with the relative improvement compared to the baseline. The best result is highlighted in bold.
Backbone
Algorithm
MRPC CoLA RTE STSB SST2 Avg
BERT-B baseline 88.45
$\pm$0.40
60.67
$\pm$0.54
68.95
$\pm$0.69
87.87
$\pm$0.12
91.97
$\pm$0.19
79.582
noisy tuning
88.43
$\pm$0.12
61.79
$\pm$0.16
71.36
$\pm$0.17
88.27
$\pm$0.94
92.32
$\pm$0.25
80.43(+0.85)
GMoE 88.63
$\pm$0.53
61.25
$\pm$2.36
70.28
$\pm$0.68
88.63
$\pm$0.65
92.28
$\pm$0.24
80.21(+0.63)
EMoE-learn
89.05
$\pm$0.23
62.46 $\pm$1.01 70.40 $\pm$1.28 88.47
$\pm$0.74
92.58
$\pm$0.14
80.59(+1.01)
EMoE 89.45 $\pm$0.36 61.55
$\pm$0.67
69.68
$\pm$1.02
88.71 $\pm$0.50 92.89 $\pm$0.19 80.46(+0.87)
BERT-L baseline 89.82
$\pm$1.30
65.41
$\pm$0.47
74.89
$\pm$1.39
89.87
$\pm$0.28
93.50
$\pm$0.24
82.70
noisy tuning
90.42
$\pm$0.35
64.75
$\pm$1.31
73.41
$\pm$1.62
90.05
$\pm$0.46
93.65
$\pm$0.11
82.46(-0.24)
GMoE 91.24 $\pm$0.25 64.90
$\pm$1.26
74.24
$\pm$1.04
90.00
$\pm$0.64
93.58
$\pm$0.25
82.79(+0.09)
EMoE-learn
90.57
$\pm$0.43
65.51
$\pm$0.32
74.72
$\pm$2.13
90.22
$\pm$0.49
93.73 $\pm$0.35 82.95(+0.25)
EMoE 90.74
$\pm$0.65
65.79 $\pm$1.16 76.17 $\pm$0.00 90.31 $\pm$0.43 93.58
$\pm$0.32
83.32(+0.62)
GPT2 baseline 84.46
$\pm$0.51
47.07
$\pm$1.60
67.15
$\pm$0.51
86.29
$\pm$0.29
92.13
$\pm$0.30
75.42
noisy tuning
84.15
$\pm$0.92
46.16
$\pm$2.79
67.51
$\pm$0.78
86.09
$\pm$0.38
92.13
$\pm$0.27
75.21(-0.21)
GMoE 85.07
$\pm$0.45
47.77
$\pm$3.20
67.51
$\pm$0.51
86.57
$\pm$0.29
92.35
$\pm$0.35
75.85(+0.43)
EMoE-learn
85.73 $\pm$0.09 47.24
$\pm$1.48
67.99
$\pm$0.17
86.66 $\pm$0.32 92.35
$\pm$0.38
75.99(+0.57)
EMoE 85.40
$\pm$0.77
48.00 $\pm$1.50 68.95 $\pm$0.29 86.64
$\pm$0.16
92.70 $\pm$0.22 76.34(+0.92)
Table 10: Results on various algorithms with different models and tasks. All the reported results are obtained from 3 independent experiments. OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization. The best result is highlighted in bold, and the second is marked with underline.
Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
OOD
BERT-Large (340 Million Parameters) Results
LoRA 89.97
$\pm$0.40
63.40
$\pm$0.62
72.92
$\pm$1.64
90.51 $\pm$0.18 93.16
$\pm$0.19
92.20
$\pm$0.13
87.21
$\pm$0.60
85.40
$\pm$0.07
84.35 4.86
Block 89.34
$\pm$0.84
62.10
$\pm$0.91
71.96
$\pm$1.68
90.39
$\pm$0.14
93.35
$\pm$0.43
92.04
$\pm$0.16
88.45 $\pm$0.07 86.20 $\pm$0.10 84.23(-0.12) 4.95
Block+GMoE 89.45
$\pm$0.72
63.80
$\pm$0.71
72.56
$\pm$0.29
90.29
$\pm$0.07
93.85 $\pm$0.11 92.32
$\pm$0.14
87.99
$\pm$0.06
85.92
$\pm$0.13
84.52(+0.18) 4.04
Block+EMoE-learn
89.79
$\pm$0.23
64.16
$\pm$0.87
73.16
$\pm$1.04
90.27
$\pm$0.03
93.85 $\pm$0.11 92.40 $\pm$0.06 88.01
$\pm$0.12
85.76
$\pm$0.19
84.68(+0.33) 3.94
Block+EMoE 89.77
$\pm$0.46
63.25
$\pm$0.50
71.60
$\pm$0.68
90.31
$\pm$0.09
93.69
$\pm$0.32
92.09
$\pm$0.13
88.08
$\pm$0.19
86.21 $\pm$0.16 84.38(+0.03) 5.89
EMoE 90.85 $\pm$0.61 65.33 $\pm$0.40 75.21 $\pm$1.62 90.43
$\pm$0.06
93.50
$\pm$0.33
92.23
$\pm$0.10
87.74
$\pm$0.10
85.43
$\pm$0.10
85.09(+0.74) 4.37
EMoE+LN 90.47 $\pm$0.33 64.39 $\pm$0.31 73.41 $\pm$1.04 90.54 $\pm$0.03 93.00
$\pm$0.16
92.31
$\pm$0.05
88.79 $\pm$0.17 85.50
$\pm$0.10
84.80(+0.46) 4.53
EMoE-learn 89.87
$\pm$0.50
64.00
$\pm$0.57
71.36
$\pm$1.39
90.48
$\pm$0.10
93.65
$\pm$0.33
92.40 $\pm$0.11 87.55
$\pm$0.14
85.62
$\pm$0.23
84.37(+0.02) 4.66
EMoE-learn+LN
89.9
$\pm$0.25
64.16
$\pm$1.16
72.44
$\pm$0.45
90.45
$\pm$0.10
93.42
$\pm$0.38
92.15
$\pm$0.10
87.70
$\pm$0.04
85.52
$\pm$0.24
84.47(0.12) 4.28
GPT2-XL (1.5 Billion Parameters) Results
LoRA 86.83
$\pm$0.87
60.88
$\pm$2.54
78.70
$\pm$0.59
89.07
$\pm$0.11
95.18
$\pm$0.28
91.84
$\pm$0.09
87.41
$\pm$1.74
86.93
$\pm$0.15
84.61 5.61
Block 86.59
$\pm$1.45
61.18
$\pm$1.74
79.78
$\pm$2.22
89.08
$\pm$0.15
95.45 $\pm$0.19 91.88
$\pm$0.05
87.71
$\pm$2.95
86.95
$\pm$0.08
84.83(+0.22) 5.13
Block+GMoE 87.02
$\pm$0.76
62.81
$\pm$1.51
79.78
$\pm$1.35
89.21
$\pm$0.20
95.41
$\pm$0.28
92.18
$\pm$0.11
89.10
$\pm$0.78
87.17 $\pm$0.20 85.34(+0.73) 4.33
Block+EMoE-learn
87.31
$\pm$1.23
62.24
$\pm$1.51
79.54
$\pm$0.17
89.33
$\pm$0.11
95.30
$\pm$0.09
92.20
$\pm$0.09
88.59
$\pm$1.68
87.06 $\pm$0.18 85.20(+0.59) 4.05
Block+EMoE 87.86
$\pm$0.98
62.88 $\pm$0.54 80.05 $\pm$0.29 89.18
$\pm$0.25
95.49 $\pm$0.39 92.10
$\pm$0.15
89.69
$\pm$0.15
86.87
$\pm$0.11
85.52(+0.91) 5.71
EMoE 87.75
$\pm$0.14
62.27
$\pm$0.93
80.02 $\pm$0.34 89.37
$\pm$0.30
95.41
$\pm$0.32
92.10
$\pm$0.15
89.58
$\pm$0.10
87.06
$\pm$0.25
85.45(+0.84) 3.88
EMoE+LN 88.05 $\pm$0.35 63.11 $\pm$0.51 79.90
$\pm$1.51
89.40
$\pm$0.22
95.18
$\pm$0.28
92.23
$\pm$0.11
89.70
$\pm$0.09
87.03
$\pm$0.14
85.58(+0.97) 4.39
EMoE-learn 87.93 $\pm$0.61 61.50
$\pm$1.09
79.90
$\pm$0.61
89.48 $\pm$0.24 95.18
$\pm$0.11
92.33 $\pm$0.093 89.71 $\pm$0.06 87.00
$\pm$0.19
85.38(+0.77) 4.40
EMoE-learn+LN
87.04
$\pm$1.11
62.64
$\pm$0.84
79.78
$\pm$0.59
89.50 $\pm$0.22 95.30
$\pm$0.50
92.31 $\pm$0.19 89.43
$\pm$0.35
87.00
$\pm$0.12
85.38(+0.77) 3.67
A.4.2 Full Tables and Figure Data Sources in Analysis
In the Analysis section 5.1, for the sake of clarity, we have transformed tabular data into graphs or retained only a subset of the results. The original and complete results corresponding to them are presented in this section. Figure 3 is summarized from Table 11 and Table 12. The OOD results in Table 3 are from 10. The ablation studies are from Table 14 and Table 15.
Table 11: ID and OOD results of BERT-L for different settings. ”LoRA-to-EMoE” refers to converting a model tuned using standard LoRA into EMoE for testing. On the other hand, ”EMoE-to-LoRA” involves merging a tuned EMoE model back into a standard dense model during testing.
Algorithm CoLA Gram MRPC Twitter RTE Hans SciTail STSB Sick Avg
LoRA 60.89
$\pm$2.55
41.77
$\pm$1.62
86.83
$\pm$0.87
75.42
$\pm$2.71
78.70
$\pm$1.02
60.37
$\pm$1.31
77.36
$\pm$0.73
89.22
$\pm$0.13
78.48
$\pm$0.33
72.12
LoRA-to-EMoE
61.31
$\pm$2.14
41.99
$\pm$1.56
86.83
$\pm$0.91
75.15
$\pm$3.02
78.58
$\pm$0.95
60.39
$\pm$1.32
77.24
$\pm$0.68
89.23
$\pm$0.13
78.53
$\pm$0.35
72.14
EMoE 62.69
$\pm$0.91
42.95
$\pm$0.95
87.82
$\pm$0.17
76.07
$\pm$2.12
79.54
$\pm$0.45
61.56
$\pm$1.65
78.09
$\pm$0.56
89.39
$\pm$0.31
78.57
$\pm$0.67
72.96
EMoE-to-LoRA
62.69
$\pm$0.91
42.94
$\pm$0.95
87.82
$\pm$0.17
76.06
$\pm$2.12
79.54
$\pm$0.45
61.56
$\pm$1.65
78.07
$\pm$0.58
89.39
$\pm$0.3
78.58
$\pm$0.67
72.96
Table 12: ID results of BERT-L for different settings. ”Cluster-top” refers to EMoE utilizing $\operatorname{avg-k}$ gating. ”Cluster-not-top” represents a scenario where, during gating, the top-k experts are removed. Similarly, ”Cluster-bottom” involves selecting the bottom-k experts with the lowest scores during gating. ”Random” denotes the approach of randomly selecting key values to construct experts. The terms ”top,” ”not-top,” and ”bottom” have the same meanings as in the cluster situations.
Algorithm MRPC CoLA RTE STSB SST2 QNLI Avg
LoRA 89.97
$\pm$0.40
63.40
$\pm$0.62
72.92
$\pm$1.64
90.51
$\pm$0.18
93.16
$\pm$0.19
92.20
$\pm$0.13
83.69
Cluster-top 90.85
$\pm$0.61
65.33
$\pm$0.40
75.21
$\pm$1.62
90.54
$\pm$0.03
93.50
$\pm$0.33
92.23
$\pm$0.10
84.61(+0.92)
Cluster-not-top 89.61
$\pm$0.76
63.21
$\pm$0.44
72.56
$\pm$1.28
90.31
$\pm$0.07
93.12
$\pm$0.34
92.14
$\pm$0.18
83.49(-0.20)
Cluster-bottom 89.21
$\pm$0.69
63.08
$\pm$1.09
71.72
$\pm$0.34
90.15
$\pm$0.18
92.97
$\pm$0.19
92.13
$\pm$0.31
83.21(-0.48)
Random-top 89.88
$\pm$0.75
63.26
$\pm$0.39
72.56
$\pm$1.06
90.33
$\pm$0.05
93.35
$\pm$0.00
92.14
$\pm$0.20
83.59(-0.11)
Random-not-top 90.09
$\pm$0.75
63.35
$\pm$0.34
72.44
$\pm$1.19
90.44
$\pm$0.07
93.31
$\pm$0.25
92.20
$\pm$0.16
83.64(-0.05)
Random-bottom 89.47
$\pm$0.23
63.17
$\pm$0.98
71.96
$\pm$0.74
90.30
$\pm$0.23
93.11
$\pm$0.25
92.10
$\pm$0.11
83.35(-0.34)
Table 13: Raw OOD performances across 13 tasks. Average results with standard deviation and best results are reported separately. Due to the large deviation across seeds overall methods, we use Friedman rank metrics (Friedman, 1940).
Algorithm
Twitter-M
GrammarTest
Hans SciTail Sick-S
NewsQA
Amazon Flipkart Imdb Yelp MRPC
Twitter-Q
Sick-M
BERT-Large (340 Million Parameters) Results (mean and standard deviation)
LoRA 80.09
$\pm$0.49
43.55
$\pm$1.34
55.51
$\pm$1.94
81.13
$\pm$0.88
81.63
$\pm$0.16
78.22
$\pm$0.17
89.06
$\pm$1.81
91.79
$\pm$0.27
84.97
$\pm$1.32
88.40
$\pm$2.06
71.73
$\pm$0.46
78.10
$\pm$0.33
53.17
$\pm$0.92
LoRA+block 80.70
$\pm$0.35
42.94
$\pm$0.47
54.35
$\pm$1.08
80.99
$\pm$0.33
81.58
$\pm$0.18
78.33
$\pm$0.24
88.21
$\pm$1.20
91.26
$\pm$0.96
84.10
$\pm$0.85
89.13
$\pm$0.20
71.57
$\pm$0.87
79.22
$\pm$0.59
53.67
$\pm$0.77
GMoE 81.17
$\pm$0.54
46.76
$\pm$1.91
58.73
$\pm$0.74
79.41
$\pm$0.10
81.35
$\pm$0.34
78.05
$\pm$0.28
89.27
$\pm$1.38
91.77
$\pm$0.07
84.93
$\pm$1.09
88.70
$\pm$1.77
71.08
$\pm$0.69
78.70
$\pm$0.37
53.85
$\pm$1.27
Block+EMoE-learn
81.09
$\pm$0.47
43.29
$\pm$1.42
57.55
$\pm$3.48
77.72
$\pm$0.61
80.78
$\pm$0.49
78.32
$\pm$0.13
90.17
$\pm$0.32
91.82
$\pm$0.18
85.71
$\pm$0.19
89.62
$\pm$0.24
72.71
$\pm$0.23
77.85
$\pm$1.03
54.69
$\pm$0.15
Block+EMoE 81.07
$\pm$0.53
46.33
$\pm$1.40
53.50
$\pm$0.97
79.62
$\pm$0.45
80.15
$\pm$0.61
79.03
$\pm$0.25
88.81
$\pm$0.19
91.33
$\pm$0.23
84.92
$\pm$0.07
88.87
$\pm$0.26
70.59
$\pm$1.00
74.62
$\pm$2.70
53.04
$\pm$1.82
EMoE 80.82
$\pm$0.10
44.18
$\pm$0.63
61.30
$\pm$2.24
76.93
$\pm$2.52
81.19
$\pm$0.27
77.87
$\pm$0.19
90.25
$\pm$0.10
92.14
$\pm$0.25
85.63
$\pm$0.30
90.21
$\pm$0.14
71.24
$\pm$1.22
78.00
$\pm$0.87
51.55
$\pm$0.88
EMoE+ln 81.04
$\pm$0.10
45.05
$\pm$0.88
58.17
$\pm$1.81
77.92
$\pm$1.64
81.27
$\pm$0.17
77.40
$\pm$0.25
90.08
$\pm$0.07
91.64
$\pm$0.54
85.67
$\pm$0.29
89.23
$\pm$0.67
72.14
$\pm$0.76
77.34
$\pm$1.33
52.81
$\pm$0.48
EMoE-learn 81.08
$\pm$0.24
46.09
$\pm$2.12
57.56
$\pm$0.43
77.52
$\pm$2.15
81.25
$\pm$0.13
78.16
$\pm$0.30
89.84
$\pm$0.13
91.87
$\pm$0.28
84.97
$\pm$0.33
89.76
$\pm$0.33
71.32
$\pm$2.03
77.47
$\pm$0.76
53.29
$\pm$0.45
EMoE-learn+ln
81.34
$\pm$0.30
44.46
$\pm$1.67
58.38
$\pm$0.59
77.78
$\pm$1.71
81.13
$\pm$0.26
78.28
$\pm$0.13
90.02
$\pm$0.29
91.48
$\pm$0.30
85.62
$\pm$0.46
89.87
$\pm$0.30
71.90
$\pm$0.61
77.38
$\pm$1.77
53.44
$\pm$0.47
BERT-Large (340 Million Parameters) Results (best result)
LoRA 80.75 45.42 58.24 82.03 81.79 78.46 90.35 92.05 85.95 90.37 72.06 78.41 54.44
LoRA+block 81.19 43.56 55.48 81.45 81.83 78.63 89.49 92.48 85.25 89.17 72.79 79.88 54.71
GMoE 81.78 49.14 59.72 79.55 81.66 78.41 90.31 91.86 85.93 90.42 72.06 79.02 55.16
Block+EMoE-learn
81.72 44.45 61.70 78.47 81.38 78.46 90.59 92.04 85.93 89.95 73.04 78.89 54.91
Block+EMoE 81.49 47.77 54.57 80.15 80.85 79.32 88.98 91.66 85.00 89.23 71.81 77.14 55.56
EMoE 80.95 45.05 64.31 79.64 81.55 78.10 90.37 92.42 86.03 90.34 72.55 79.23 52.24
EMoE+ln 81.17 45.71 60.73 79.84 81.48 77.63 90.15 92.33 86.01 90.60 72.79 79.53 53.28
EMoE-learn 81.34 48.75 58.11 79.89 81.25 78.58 89.93 92.23 85.38 90.21 74.02 78.33 53.91
EMoE-learn+ln
81.72 46.82 59.09 79.60 81.41 78.47 90.27 91.72 86.11 90.17 72.55 78.67 54.10
GPT2-XL (1.5 Billion Parameters) Results (mean and standard deviation)
LoRA 75.42
$\pm$2.71
41.78
$\pm$1.62
60.37
$\pm$1.32
77.36
$\pm$0.73
78.48
$\pm$0.33
78.57
$\pm$0.59
89.91
$\pm$0.73
90.05
$\pm$0.44
85.57
$\pm$1.34
88.85
$\pm$0.38
68.22
$\pm$1.27
73.60
$\pm$2.87
57.39
$\pm$0.34
LoRA+block 76.87
$\pm$2.47
41.09
$\pm$1.80
58.67
$\pm$0.97
78.57
$\pm$1.23
77.66
$\pm$0.25
78.78
$\pm$0.55
90.11
$\pm$0.71
91.54
$\pm$0.94
83.54
$\pm$0.57
88.85
$\pm$1.01
69.53
$\pm$0.64
68.22
$\pm$7.84
58.53
$\pm$0.46
GMoE 75.28
$\pm$3.67
42.50
$\pm$1.37
62.59
$\pm$0.37
77.89
$\pm$0.56
78.33
$\pm$0.58
79.18
$\pm$0.04
90.11
$\pm$0.38
88.64
$\pm$4.52
83.23
$\pm$0.40
89.47
$\pm$0.26
70.10
$\pm$1.39
74.61
$\pm$3.71
58.29
$\pm$0.78
Block+EMoE-learn
74.80
$\pm$5.15
44.15
$\pm$1.95
61.67
$\pm$1.18
78.18
$pm$1.04
78.27
$\pm$0.21
78.97
$\pm$0.20
90.22
$\pm$0.59
91.04
$\pm$0.65
83.82
$\pm$0.54
89.49
$\pm$0.67
69.93
$\pm$0.50
73.41
$\pm$3.94
57.58
$\pm$0.57
Block+EMoE 75.37
$\pm$3.04
39.33
$\pm$2.21
58.32
$\pm$2.40
72.05
$\pm$3.04
77.70
$\pm$0.59
79.32
$\pm$0.06
90.24
$\pm$0.36
91.79
$\pm$0.27
83.79
$\pm$0.27
89.23
$\pm$0.38
63.24
$\pm$8.56
70.60
$\pm$2.89
56.85
$\pm$0.57
EMoE 76.07
$\pm$2.12
42.95
$\pm$0.95
61.56
$\pm$1.65
78.09
$\pm$0.56
78.57
$\pm$0.67
78.87
$\pm$0.38
90.39
$\pm$0.55
91.87
$\pm$0.22
83.55
$\pm$1.04
89.14
$\pm$1.39
67.73
$\pm$1.55
74.06
$\pm$4.79
57.33
$\pm$0.59
EMoE+ln 74.35
$\pm$3.50
42.24
$\pm$0.89
61.88
$\pm$2.45
78.32
$\pm$0.95
78.67
$\pm$0.62
79.02
$\pm$0.30
89.51
$\pm$0.90
91.19
$\pm$0.56
83.12
$\pm$1.90
89.29
$\pm$1.09
69.20
$\pm$1.03
72.74
$\pm$5.10
57.62
$\pm$0.02
EMoE-learn 74.61
$\pm$4.75
40.50
$\pm$0.46
59.50
$\pm$3.74
80.04
$\pm$0.46
78.27
$\pm$0.59
79.06
$\pm$0.50
90.59
$\pm$0.66
91.69
$\pm$0.59
83.23
$\pm$1.54
89.79
$\pm$1.35
67.48
$\pm$1.30
73.62
$\pm$1.12
56.22
$\pm$0.15
EMoE-learn+ln
77.17
$\pm$1.58
42.52
$\pm$0.41
58.90
$\pm$4.16
79.56
$\pm$0.81
78.27
$\pm$0.54
79.19
$\pm$0.32
90.23
$\pm$0.86
91.98
$\pm$0.35
83.31
$\pm$1.08
88.50
$\pm$1.92
68.63
$\pm$1.51
71.26
$\pm$3.49
56.69
$\pm$0.37
GPT2-XL (1.5 Billion Parameters) Results (best result)
LoRA 78.17 43.93 61.77 78.39 78.87 79.12 90.52 91.03 84.81 89.39 69.12 77.48 57.79
LoRA+block 80.17 43.09 59.52 79.45 78.33 79.42 90.94 92.87 84.12 90.28 70.10 74.55 57.64
GMoE 80.23 43.96 63.05 78.68 78.00 79.33 90.57 92.27 83.69 89.81 71.08 79.79 59.36
Block+EMoE-learn
80.04 46.01 62.74 79.65 79.06 79.24 90.67 91.94 84.45 90.42 70.59 78.78 58.28
Block+EMoE 79.02 41.55 61.64 76.30 78.33 79.40 90.68 92.00 84.12 89.60 73.31 73.31 57.64
EMoE 77.58 43.84 63.78 78.75 79.29 79.26 91.17 92.19 84.42 90.22 69.85 80.81 58.01
EMoE+ln 77.92 43.39 65.32 79.16 79.25 79.35 90.48 91.74 85.73 90.79 70.59 79.81 57.87
EMoE-learn 79.35 41.15 64.56 80.40 78.76 79.76 91.32 92.47 84.49 90.77 69.12 74.77 56.42
EMoE-learn+ln
79.35 43.01 64.32 80.57 79.02 79.63 91.40 92.41 84.84 90.57 70.34 74.77 57.13
Table 14: Results of different MoEs Configurations
Algorithm MRPC CoLA RTE STSB Avg
Dense 86.83
$\pm$0.87
60.88
$\pm$2.54
78.70
$\pm$0.59
89.07
$\pm$0.11
78.87
EMoE 88.05
$\pm$0.35
63.11
$\pm$0.21
80.02
$\pm$0.34
89.37
$\pm$0.24
80.14
EMoE-learn 87.93
$\pm$0.61
62.87
$\pm$0.71
79.90
$\pm$0.61
89.40
$\pm$0.08
80.03
EMoE-last-every2
87.27
$\pm$0.47
61.60
$\pm$0.63
79.18
$\pm$0.17
89.38
$\pm$0.24
79.36
EMoE-learn-learn-every2
87.26
$\pm$0.21
61.82
$\pm$1.10
78.46
$\pm$1.22
89.31
$\pm$0.15
79.21
EMoE-every2
86.78
$\pm$0.34
59.21
$\pm$0.79
77.38
$\pm$1.04
89.31
$\pm$0.06
78.17
EMoE-learn-every2
86.71
$\pm$1.32
54.02
$\pm$0.47
74.25
$\pm$0.74
88.51
$\pm$0.31
75.87
Table 15: Comparison of EMoE and Dense Results with Different Training Data Fraction
Data CoLA MRPC RTE STSB SST2 QNLI Average
Fraction EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense Diff
1.0 62.27 60.88 87.75 86.83 80.02 78.70 89.37 89.07 95.41 95.18 92.10 91.84 0.74
0.9 61.58 60.01 87.52 86.49 79.87 77.85 89.18 89.07 95.41 95.16 92.05 91.94 0.85
0.8 60.89 59.28 86.58 86.35 79.22 76.77 88.98 88.99 95.34 95.06 91.94 91.83 0.78
0.7 59.29 58.25 86.10 85.56 77.98 76.77 87.95 87.95 95.41 95.06 92.04 91.74 0.57
0.6 58.91 57.93 85.61 84.76 76.29 75.45 86.70 86.17 95.26 95.03 91.83 91.54 0.62
0.5 55.18 53.89 84.91 84.76 76.53 75.21 85.63 85.59 95.19 94.91 91.23 91.12 0.53
0.3 50.17 50.29 82.80 82.59 73.52 72.44 80.23 79.08 94.82 94.72 90.45 90.26 0.44
0.1 46.47 45.54 78.17 77.85 63.05 63.17 62.33 60.81 94.49 94.38 88.39 88.32 0.47
A.5 More Visualization Results
In this part, we demonstrate more gating visualization results on SST-2, STSB, MRPC, and RTE in Figure 8. These results are consistent with earlier findings: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating. This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022).
Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from dense vs repetitions of dense). The vertical axis illustrates training steps, with top-down arrangement signifying begin-end; the horizontal axis represents expert selection frequency within 1K steps. (a), (b) and (c) correspond respectively to EMoE, EMoE-learn, and GMoE configurations. The subplots from top to bottom are results for SST-2, STS-B, MRPC, and RTE.

### v2

@@
-Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?
+Unlocking Emergent Modularity in Large Language Models
 Abstract
-Incorporating modular designs into neural networks demonstrates superior out-of-generalization, learning efficiency, etc.
-Existing modular neural networks are generally explicit because their modular architectures are pre-defined, and individual modules are expected to implement distinct functions.
-Conversely, recent works reveal that there exist implicit modular structures in standard pre-trained transformers, namely Emergent Modularity.
-They indicate that such modular structures exhibit during the early pre-training phase and are totally spontaneous.
-However, most transformers are still treated as monolithic models with their modular natures underutilized.
-Therefore, given the excellent properties of explicit modular architecture, we explore whether and how dense pre-trained transformers can benefit from emergent modular structures.
-To study this question, we construct Emergent Mixture-of-Experts (EMoE).
-Without introducing additional parameters, EMoE can be seen as the modular counterpart of the original model and can be effortlessly incorporated into downstream tuning.
-Extensive experiments (we tune 1785 models) on various downstream tasks (vision and language) and models (22M to1.5B) demonstrate that EMoE effectively boosts in-domain and out-of-domain generalization abilities.
-Further analysis and ablation study suggest that EMoE mitigates negative knowledge transfer and is robust to various configurations.
-Code is available at https://github.com/qiuzh20/EMoE
+Modular Neural Networks (MNNs) demonstrate various advantages over monolithic models.
+Existing MNNs are generally explicit: their modular architectures are pre-defined, with individual modules expected to implement distinct functions.
+Recent works reveal that there exists implicit modularity in standard pre-trained transformers, namely Emergent Modularity.
+They indicate that such modular structures spontaneously exhibit during the early pre-training phase.
+Despite the benefits of modularity, most Language Models (LMs) are still treated as monolithic models in the pre-train and fine-tune paradigm, with their emergent modularity locked and underutilized.
+In this work, focusing on unlocking the emergent modularity in LMs, we showcase that standard LMs could be fine-tuned as their Mixture-of-Expert (MoEs) counterparts without introducing any extra parameters.
+Such MoEs are derived from emergent modularity and are referred to as Emergent MoEs (EMoE).
+Our experiments demonstrate that fine-tuning EMoE effectively improves downstream in-domain and out-of-domain generalization compared with vanilla fine-tuning.
+Our analysis and ablation studies further illustrate that it is robust to various configurations and can scale up to Large Language Models (i.e., Llama2-7B and Llama-30B).
+Code is available at this repo.
+Unlocking Emergent Modularity in Large Language Models
+12Zihan Qiuthanks: Equal contributionthanks: Work done while interning at HKUST.
+, 3Zeyu Huang, 1Jie Futhanks: Corresponding author 1CSE, HKUST     2IIIS, Tsinghua University, 3ILCC, University of Edinburgh qzh11628@gmail.com, zeyu.huang@ed.ac.uk, jiefu@ust.hk
 1 Introduction
-Modularity attracts considerable attention from the artificial intelligence community (Auda & Kamel, 1999).
-Neural networks with modular designs, termed Modular Neural Networks (MNN), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), out-of-distribution (OOD) generalization abilities (Goyal & Bengio, 2020; Weiss et al., 2022), and data efficiency (Bengio et al., 2020).
+Modularity attracts considerable attention from the artificial intelligence community (Auda and Kamel, 1999).
+Neural networks with modular designs, termed Modular Neural Networks (MNNs), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), data efficiency (Bengio et al., 2020), and better generalization abilities (Goyal and Bengio, 2020; Weiss et al., 2022).
 Typical MNNs are usually explicitly modular.
-Their modular structure is pre-defined and they are expected to achieve a divide-and-conquer solution for the given task.
-For example, Andreas et al.
-(2016) design separate visual and language functional modules and jointly train them to solve visual question-answering tasks.
+They have a pre-defined modular structure and are expected to achieve a divide-and-conquer solution for the given task.
 Among various MNNs, Mixture-of-Experts (MoEs) employ a conditional computation strategy where different submodules - so-called experts - are expected to be activated by different types of inputs.
-MoEs sees substantial success in various domains (Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and therefore becomes a widespread modular neural architecture.
-Apart from developing explicit MNN,  Csordás et al.
-(2021); Agarwala et al.
-(2021) study whether standard neural networks become implicitly modular after training and discover spontaneously emerged modular structure in small-scale CNNs and LSTMs.
-For more complicated large-scale pre-trained transformers, initial observations (Zhang et al., 2022b; Li et al., 2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs).
-Specifically, they find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process.
-However, sparsity does not imply modularity.
-Sparsely activated FFNs neurons do not ensure to have remarkable functional division.
-Therefore,  Zhang et al.
-(2023) further utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs.
-They observe a strong correlation between activation and specific tasks, further discovering clear function-based neuron grouping of the model and summarizing such phenomenon as Emergent Modularity.
-Figure 1: (a) Existing literatures (Geva et al., 2021; 2022) suggest that the FFNs in transformers can be viewed as key-value memories.
+MoEs see substantial success in various domains (Fedus et al., 2022; Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and they are therefore a widespread modular neural architecture.
+Apart from explicit MNNs, some research finds that modular structures spontaneously emerge during training, not only in small-scale CNNs or LSTMs (Csordás et al., 2021; Agarwala et al., 2021), but also large-scale pre-trained transformer models.
+Zhang et al.
+(2022b); Li et al.
+(2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs) in pre-trained transformer models.
+They find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process, respectively.
+Furthermore,  Zhang et al.
+(2023) utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs.
+They observe a strong correlation between neuron activation and specific tasks, further discovering clear function-based neuron grouping of the pre-trained T5 model (neurons with similar functions are usually co-activated).
+They summarize such phenomenon as Emergent Modularity (EM).
+Though modularity emerges, pre-trained language models are generally treated as monolithic models in the standard pre-train and fine-tune paradigm.
+It is natural to ask whether their EM and the potential improvements brought by EM are locked in this process.
+In this paper, we advocate unlocking the EM in the pre-trained language models could bring generalization improvements for downstream tasks.
+Specifically, we split certain FFNs layers of the original model into MoEs layers.
+The MoEs is derived according to the EM in that layer and can be regarded as the externalization of the EM.
+Hence, the obtained MoEs model is called Emergent MoEs (EMoE).
+We then fine-tune the obtained EMoE model to investigate whether unlocking EM encourages downstream task performance.
+We validate our empirical findings with various models, evaluation benchmarks, fine-tuning methods (parameter-efficient tuning and full fine-tuning).
+We find that fine-tuning EMoE achieves stronger generalization performance than vanilla fine-tuning across various experimental settings, demonstrating that unlocking the EM of LMs boosts the models’ downstream generalization abilities.
+We provide a comprehensive analysis for EMoE: 1) We first validate that EMoE indeed unlocks EM in pre-trained language models by showcasing its task-specific expert choice.
+2) We then reveal that EMoE ameliorates the parameter updating during fine-tuning and can even be abandoned afterward.
+We want to highlight that this property further improves the practicality of EMoE as the model architecture does not change before and after fine-tuning.
+Meanwhile, our ablation studies show the EMoE’s robustness to various hyper-parameter configurations.
+3) We also conclude that EMoE could mask neurons with negative transfer effects.
+We hope our research discoveries can bring novel insights and serve as an example attempt towards further unlocking the EM of LLMs.
+2 Methodology
+2.1 Preliminaries
+Transformer FFNs are Key-Value Memories.
+The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{h\times d}$, $\mathbf{V}\in\mathbb{R}^{d\times h}$, where $h$ is embedding size and $d$ is the dimension of the hidden layer (usually $d=4h$), and a non-linear activation function $\sigma(\cdot)$.
+For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as Eq.
+1:
+$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K})\cdot\mathbf{V}.$ (1)
+More precisely, for each column $\mathbf{K}_{:,i}$ and row $\mathbf{V}_{i,:}$, Eq.1 can be rewritten as:
+$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K}_{:,i})\cdot\mathbf{V}_{i,:}$ (2)
+Following Geva et al.
+(2021, 2022); Huang et al.
+(2023), we regard columns in $\mathbf{K}$ as key vectors and rows in $\mathbf{V}$ as value vectors, the output of an FFNs network can be viewed as a weighted sum of value vectors based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K})$.
+For the rest of this paper, we refer to one key-value memory pair using neuron and the co-activation of neurons using modularity.
+Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module.
+Supposing there are $N$ experts: $\{\operatorname{FFN}^{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$.
+Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function.
+Given $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ of can be summarized as the weighted sum of the output from all experts:
+$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}^{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ (3)
+When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation, termed sparse MoEs.
+2.2 Emergent Mixture-of-Expert
+According to Eq.
+2 and Eq.
+3, we find that FFNs internally resemble MoEs if we consider keys as the gating module and values as the group of experts, which inspire us to transform existing FFNs into sparse MoEs to unlock its modular potentials.
+Since our research goals mainly focus on EM, a preferred approach to externalizing EM into sparse MoEs should not introduce additional parameters, training, and data, which may result in impractical or undesired biases.
+Therefore, after splitting original neurons into different groups to construct different experts, each group’s average of key vectors is calculated to form the gating module.
+Details are described and illustrated in Fig.
+1.
+Figure 1: (a) Existing literatures (Geva et al., 2021, 2022) suggest that the FFNs in transformers can be viewed as key-value memories.
 They regarded the input as a query, the first layer as keys, and the second as values.
@@
 (2) The FFNs block can be partitioned into experts by clustering keys.
-(3) Afterward, experts’ keys averages are used as the gating weights.
+(3) Afterward, experts’ key averages are used as the gating weights.
 The inner product between x and gating weights are used to select experts.
-However, contemporary transformers are still employed as monolithic models, with their emergent modularity underutilized.
-Motivated by the advantages of explicit modularity, we explore whether and how standard pre-trained transformers could benefit from emergent modular structures.
-Technically, we externalize the emergent modularity into the explicit MoEs, i.e., Emergent Mixcture-of-Experts (EMoE).
-As shown in Figure 1, we first (1) utilize a clustering-based method to decompose the FFNs into experts and then (2) use the Average-Key-Gating ($\operatorname{avg-k}$ gating) to route the input to corresponding experts.
-Both two steps do not require extra training processes and data.
-Then we fine-tune EMoE on a wide range of downstream tasks to investigate if standard pre-trained transformers can benefit from its ermgent modular structure.
-We evaluate EMoE following different configurations: (1) tasks from various modalities: vision (Domainbed (Gulrajani & Lopez-Paz, 2021)) and language (GLUE (Wang et al., 2019) and GLUE-X (Yang et al., 2023)); (2) different pre-trained dense transformers: ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), and GPT2 (Radford et al., 2019), ranging from 22M to 1.5B parameters.
-(3) different evaluation settings: in-domain (ID) evaluation and OOD evaluation; (4) different fine-tuning methods: full fine-tuning and parameter-efficient tuning (Hu et al., 2022).
-In total, we fine-tune 1785 models.
-Refer to caption Refer to caption
-Figure 2: EMoE achieves stronger ID and OOD performances compared with baselines in both full-fine-tuning (Left, average accuracy) and LoRA tuning settings (Hu et al., 2022) (Right, average accuracy for ID and Friedman rank (Friedman, 1940) for OOD).
-Results are summarized from Section 4.
-Our results are summarized in Figure 2.
-We find that fine-tuning EMoE achieves significant improvement compared with the vanilla fine-tuning and competitive performance with strong baseline GMoE (Li et al., 2023), demonstrating that dense pre-trained transformer can indeed benefit from emergent modular structures to achieve better generalization performance.
-We further find that EMoE can deactivate neurons with negative transfer effects, thus improving downstream performance.
-Meanwhile, our ablation studies indicate that EMoE’s insensitivity to various hyper-parameter configurations and higher data efficiency of up to 20%.
-2 Related Work
-Modular neural networks has various advantages, including interpretability (Pfeiffer et al., 2023), scalability (Chowdhery et al., 2022), multi-task learning abilities (Chen et al., 2023a), and OOD generalization abilities (Goyal et al., 2021; Li et al., 2023).
-MoEs (Szymanski & Lemmon, 1993) is currently regarded as a standard framework for developing modular neural networks (Shen et al., 2023b).
-Most MoEs models are developed by training explicit MoEs architecture from random intialization (Shazeer et al., 2017; Fedus et al., 2022b; Zhang et al., 2022a), which usually causes load imbalance issues (Zoph et al., 2022; Fedus et al., 2022a; Chi et al., 2022).
-Some recent works employ already-trained dense models as a good initialization (Nie et al., 2021).
-For example, GMoE (Li et al., 2023) and Upcycling Komatsuzaki et al.
-(2023) copy the FFNs from a trained dense network to continue pre-training or transfer learning.
-EMoE is also based on pre-trained dense models.
-However, unlike existing methods that often significantly increase in parameters (GMoE increases 50% parameters) or tend to be less effective (Upcycling extends T-5 Base (248M) to 2B parameters but still underperforms the T5-Large (783M)), EMoE achieves significant improvements without introducing extra training processes or parameters.
-This is primarily attributed to the heterogeneity of its experts and gating mechanisms, as we demonstrate in Section  5 and 6.
-In another research line,  Hod et al.
-(2021); Csordás et al.
-(2021) explore modular structures in CNNs and LSTMs.
-Some works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated dense pre-trained transformers.
-Based on their observations,  Zhang et al.
-(2023) recently explore modularity in pre-trained transformer FFNs.
-In this work, we adopt clustering-based expert construction and $\operatorname{avg-k}$ gating from MoEfication (Zhang et al., 2022b), while our motivation differs.
-Seeking to preserve the original functionality with fewer activated neurons, they split every layer in the dense model into MoEs.
-However, the results are worse than the dense model by more than ten points after splitting (Table 2 in their original paper).
-This is because the clustering-based decomposition and $\operatorname{avg-k}$ gating cannot precisely select the activated neurons.
-Without any tuning, these errors accumulate with MoEs layers and lead to severe performance dropping.
-Therefore, Zhang et al.
-(2022b) mainly focus on more complicated learning-based construction methods.
-Differently, we treat these decomposed units as heterogeneous modules to exploit their modular properties in the framework of MoEs, and observe significant improvements over monolithic models.
-3 Methodology
-3.1 Preliminaries
-Transformer Feed-Forward Networks The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{d\times h}$, $\mathbf{V}\in\mathbb{R}^{h\times d}$, and a non-linear activation function $\sigma(\cdot)$.
-For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as follows:
-$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}.$ (1)
-More precisely, for each row $\mathbf{K}_{i,:}$ and column $\mathbf{V_{:,i}}$, the Equation 1 can be rewritten as:
-$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K_{i,:}})\cdot\mathbf{V_{:,i}}$ (2)
-Regarding each row in $\mathbf{K}$ as a key vector and each column in $\mathbf{V}$ as the value vector, the output of an FFNs network can be viewed as a weighted sum of values based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K}^{T})$.
-Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module.
-Supposing there are $N$ experts: $\{\operatorname{FFN}_{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$.
-Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function.
-Given an input $\mathbf{x}\in\mathbb{R}^{h}$, let $\mathbf{y}_{n}=\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ be the output of the $n$-th expert, and then the output $\mathbf{y}\in\mathbb{R}^{h}$ of can also be summarized as the weighted sum of the output from all experts:
-$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)y_{n},$ (3)
-When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation.
-3.2 Emergent Mixture-Of-Expert
-The ideal EMoE should reflect the implicit modular nature of dense pre-trained transformers using the form of MoEs.
-A preferred approach to constructing EMoE should avoid introducing additional parameters, training, and data, which may result in undesired biases and are impractical.
-Although more sophisticated constructing methods can be proposed, we adopt a method from MoEfication Zhang et al.
-(2022b).
-The method contains two steps: (1) clustering-based expert construction and (2) $\operatorname{avg-k}$ gating.
-It is simple and fulfills our needs, as demonstrated in our experiments.
 Clustering-based Experts Construction.
-Since key-value pairs with similar key vectors tend to be co-activated, we split them into separate experts according to their key vectors.
-Specifically, given a trained FFNs layer $\operatorname{FFNs}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen & Fränti, 2014) to partition all key vectors $\mathbf{K}$ into $N$ experts on average, so each group contains $\frac{d}{N}$ key-value pairs (usually $d=4h$).
-Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract key-value pair $(\mathbf{K_{j,:}},\mathbf{V_{:,j}})$ to form the $i$th expert $\operatorname{FFN}(\cdot;\mathbf{K}_{i},\mathbf{V}_{i})$, as depicted in Figure 1(b).
-After that, the computation of each expert proceeds as Equation 1.
+We aim to ensure that neurons that tend to be co-activated are divided into the same group.
+Since neurons with similar key vectors tend to be co-activated according to Eq.
+2, we split them into separate experts by clustering their key vectors.
+Specifically, given a FFNs layer $\operatorname{FFN}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen and Fränti, 2014) (more details are in Appendix A.0.1) to partition all key vectors $\mathbf{K}$ into $N$ experts on average.
+Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract neuron $(\mathbf{K}_{:,j},\mathbf{V}_{j,:})$ to form the $i$-th expert $\operatorname{FFN}(\cdot;\mathbf{K}^{i},\mathbf{V}^{i})$ as depicted in Fig.
+1(b).
+After that, the computation of each expert follows Eq.
+1.
 Avg-k Gating.
-Given an input $\mathbf{x}$ and $N$ experts, a qualified gate for EMoE should route the input $\mathbf{x}$ to the experts who contribute most to the model’s output.
-We construct the gating module by averaging each expert’s keys.
-The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $n$th column $\mathbf{G}_{,:n}$ can be calculated as follows:
+As discussed, we do not want to introduce extra trainable parameters when externalizing EM.
+Therefore, we construct the gating module by averaging each expert’s keys, which should route the input $\mathbf{x}$ to the experts who tend to have larger activation scores and thus contribute more to the model’s output.
+The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $i$-th column $\mathbf{G}_{:,i}$ can be calculated as follows:
 […]
-And then the corresponding gating score for $i$-th expert is:
-$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G}^{T};k,\text{dim=1})\\ &0~{}~{}\text{ else}\end{cases}$ (5)
-where $\operatorname{Top-K}(\cdot;k,\text{dim})$ return indices of $k$ largest element of the given input along a given dimension.
-Using 0,1 score, $\operatorname{avg-k}$ gating reduces the weights’ influence in Equation 3).
-Notably, observe:
-$\mathbf{x}\cdot\mathbf{G}_{:,n}^{T}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{n},\text{dim}=0)=\mathbf{x}\cdot\frac{N}{h}\sum_{j\in E_{i}}\mathbf{K}^{n}_{j,:}=\frac{N}{h}\sum_{j}a_{j}.$ (6)
+And then the gating score for $i$-th expert is:
+$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G};k)\\ &0~{}~{}\text{ else}\end{cases}$ (5)
+where $\operatorname{Top-K}(\cdot;k)$ returns indices of $k$ largest element of the given input along.
+As the gating score is the average of activation scores (before $\sigma(\cdot)$) of neurons in that expert:
+$\mathbf{x}\cdot\mathbf{G}_{:,i}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{i})=\frac{N}{d}\sum_{j\in E_{i}}\mathbf{x}\mathbf{K}^{i}_{:,j}=\frac{N}{d}\sum_{j}a_{j}.$ (6)
 a larger value of gating score $g_{i}$ implies more activated keys within the corresponding expert.
 Consequently, the expert could potentially contribute more to the output $\mathbf{y}$ for input $\mathbf{x}$.
-During downstream tuning, gating weights are updated along with the FFNs parameters using Equation 4.
-3.3 Discussion
-Though adopting a method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios.
-MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons.
-However, the method introduces deviations from FFNs as it can mistakenly eliminate useful neurons.
-As these deviations accumulated across layers, the performances severely drop by over 10 points on original tasks.
-Therefore, they focus on more precise expert construction and gating to preserve performances.
-In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks.
-We summarize their main differences in Appendix A.1.
-As one of our most related work, GMoE (Li et al., 2023) introduce MoEs by copying the trained FFNs.
-They further demonstrate introducing MoEs every 2 layers leads to significant degradation.
-Accordingly, we only introduce EMoEs in certain layers since efficiency isn’t our primary concern.
-We also conduct ablation studies about MoEfying every 2 layers and get consistent results, more details are in Appendix A.4.2 Table 14.
+During downstream tuning, gating weights are tied with the FFNs parameters using Eq.
+4.
 3 Experiments
+Configurations: We first evaluate EMoE using BERT and GPT2 series models.
+We employ GLUE (Wang et al., 2019b) and GLUE-X (Yang et al., 2023) for benchmarking in-domain (ID) and out-of-domain (OOD) performance of the fine-tuned model, respectively.
+We mainly present the experimental results when employing LoRA (Hu et al., 2022) to fine-tune the pre-trained language models for two reasons: (1) with the increasing scale of pre-trained models, parameter-efficient tuning (Houlsby et al., 2019) can scale up to very large language models and thus becomes more practical.
+(2) standard LoRA weights are added in each self-attention block, and the parameters in FFNs will not be updated, allowing us to investigate whether leveraging EM, even without fine-tuning the parameters of EMoE, can yield improvements.
+We present EMoE full fine-tuning results in the Appendix B.1.
+Besides, we scale up EMoE on Llama2-7B and Llama-30B (Touvron et al., 2023) for further validation.
+We instruct-tuning models on Alpaca and testing them on the MMLU benchmark (Hendrycks et al., 2021).
+For more details about datasets, evaluation metrics and computation cost, please refer to the Appendix D and E.
+Table 1: ID and OOD performance of EMoE and baseline models.
+All the reported results are obtained from 3 independent experiments.
+OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for OOD generalization.
+The best result is highlighted in bold.
 […]
-Baselines On Domainbed, our baselines include (1) vanilla ViT, as it is a strong OOD baseline under fair configurations suggested by Gulrajani & Lopez-Paz (2021).
-(2) GMoE (Li et al., 2023), which is the latest state-of-the-art.
-GMoE constructs MoEs by replicating FFNs from the two-to-last and fourth-to-last transformer blocks of a pre-trained ViT and therefore is the most relevant to EMoE.
-For language tasks, besides vanilla backbone and GMoE, we implement (3) noisy tuning (Wu et al., 2022), which also improves adaptation for free by adding uniform distribution noise proportional to the standard deviation of the pre-trained weights before tuning.
-(4) EMoE-learn, an ablation method, where the gating function is learned (same as GMoE) during fine-tuning instead of employing $\operatorname{avg-k}$ gating.
+[…]
+BERT-Large (340M Parameters)
+LoRA 89.97 63.40 72.92 90.51 93.16 92.20 87.21 85.40 84.35 4.86
+GMoE 89.45 63.80 72.56 90.29 93.85 92.32 87.99 85.92 84.52(+0.18) 4.04
+EMoE-learn
+89.87 64.00 71.36 90.48 93.65 92.40 87.55 85.62 84.37(+0.02) 4.66
+EMoE 90.85 65.33 75.21 90.43 93.50 92.23 87.74 85.43 85.09(+0.74) 4.37
+GPT2-XL (1.5B Parameters)
+LoRA 86.83 60.88 78.70 89.07 95.18 91.84 87.41 86.93 84.61 5.61
+GMoE 87.02 62.81 79.78 89.21 95.41 92.18 89.10 87.17 85.34(+0.73) 4.33
+EMoE-learn
+87.93 61.50 79.90 89.48 95.18 92.33 89.71 87.00 85.38(+0.77) 4.40
+EMoE 87.75 62.27 80.02 89.37 95.41 92.10 89.58 87.06 85.45(+0.84) 3.88
+Baselines Our baselines include (1) vanilla LoRA-tuning: add LoRA weights to the q and v projections in attention layers; (2) GMoE (Li et al., 2023): Instead of splitting, GMoE replicates FFNs layer and train new gating layers to introduce MoE structure from the two-to-last and fourth-to-last transformer blocks in the original pre-trained models.
+Since GMoE copies the FFNs of the pre-trained model to form the MoEs, it is ineffective if introduced MoEs aren’t tuned.
+Thus, we conduct experiments with LoRA tuning for GMoE and tune the transformer block where the original FFNs are replaced.
+(3) EMoE-learn: an ablation method, where the gating function is learned (same as GMoE) during fine-tuning.
 This helps us better understand the effect of $\operatorname{avg-k}$ gating.
+[…]
+Evaluation Metrics All experiments except the scale-up ones are repeated three times, and the average is presented in the main section.
+Full results are in Appendix F.1.
+For the OOD metric, we follow GLUE-X (Yang et al., 2023) and employ the Friedman rank (Friedman, 1940) [^1] Our evaluation includes nine methods detailed, demonstrating the relationships among these methods.
+Therefore, these values can’t provide a direct comparison with the results in the GLUE-X paper $\text{rank}_{f}=\frac{1}{n}\sum_{i=1}^{n}\text{rank}_{i}$.
+For each method under the same backbone, $\text{rank}_{i}$ is produced based on each dataset’s best and average results.
+For the 13 OOD tasks used in GLUE-X, each method generates 26 $\text{rank}_{i}$ values.
+The OOD results presented in Table 1 represent the mean of all these $\text{rank}_{i}$ values.
+The original results of each task can be found in the Appendix F.1.
+Results with BERT and GPT2 According to Tab.
+1: (1) EMoE demonstrates enhancements compared to vanilla LoRA-tuning.
+Notably, EMoE also achieves results comparable to GMoE on BERT-large and outperforms GPT2-XL with much fewer learnable parameters.
+(2) While EMoE-learn’s results are better than EMoE in several tasks (STSB, QNLI), EMoE exhibits higher stability than EMoE-learn and delivers superior overall results.
+(3) MoEs structures improve OOD performance (GMoE, EMoE, EMoE-learn vs.
+LoRA).
+Table 2: Instruction LoRA-tuned Llama models evaluated (in accuracy) on the MMLU benchmark.
+Times are wall-clock computation times for training 3 epochs on Alpaca using the same GPU devices.
+When $N=K=1$, it represents the vanilla LoRA tuning.
+Experts (N) topk (k) MMLU score times (s) FLOPS ($10^{17}$)
+Llama2-7B
+1 1 46.96 1396 6.92
+64 16 47.58 1545 7.03
+64 32 47.37 1521 7.13
+Llama-30B
+1 1 56.18 6943 22.5
+256 64 57.11 6955 22.5
+256 128 56.64 6974 22.4
+Refer to captionRefer to caption
+Figure 2: Left: Activations of neurons (z-axis denotes activation value) in FFNs of a pre-trained transformer models.
+Middle: By clustering the keys in the FFNs layer and rearranging the activation scores accordingly, modular patterns of neuron activation emerge.
+Right: The heat map between experts and tasks.
+It is observed that the activation of experts is task-dependent, while some experts are generally shared across different tasks.
+Results with Llama Unlike GMoE, EMoE does not introduce any additional trainable parameters or procedures so that we can scale EMoE up to models of 7B and 30B sizes to validate its effectiveness further.
+From Tab.
+2: 1.
+EMoE consistently yields improvements compared with LoRA tuning with negligible additional computation.
+2.
+Because EMoE does not introduce any additional trainable parameters, the choice of K and N proportions remains large (can be even larger) and is applicable in larger-scale models.
+While the performance varies when employing different $N$ and $K$, EMoE consistently outperforms the vanilla LoRA tuning.
+This suggests that although additional hyperparameters $N$ and $K$ are introduced, they do not lead to usability challenges.
+Please refer to Sec.
+5 for more ablation results on hyperparameters.
+Additional Results Moreover, we conduct experiments with full fine-tuning under comprehensive evaluation settings: 1.
+Vision OOD benchmark Domainbed (Gulrajani and Lopez-Paz, 2021), full fine-tuning ViT-Small (22M) and ViT-Base (86M).
+2.
+GLUE benchmark, fully fine-tuning BERT-Base, BERT-Large, and GPT2-Small.
+3.
+Full fine-tuning Llama2-7B.
+We present detailed results and analyses in Appendix B.1.
+According to the results, EMoE brings consistent improvements over standard full finetuning across various tasks and model scales.
+On the vision OOD benchmark Domainbed, which strictly controls evaluation metrics, EMoE achieved results comparable to the state-of-the-art GMoE.
+Furthermore, when full-finetuned with Alpaca, EMoE exhibited a notable improvement of 1.58 on MMLU to the standard tuning baseline.
+These findings underscore the effectiveness of EMoE in enhancing model performance across various architectures and tasks.
+4 Analysis
+4.1 Does EMoE Unlock Emergent Modularity?
+We first investigate whether simple key-vector-based clustering partitioning could capture the modular pattern of the neuron activation.
+The activation scores of different neurons on different inputs are visually shown in the Fig.
+2 left and middle.
+Before clustering (Fig.
+2 left), the activation of neurons seems random.
+After rearranging those activation scores according to the EMoE partition (Fig.
+2 middle), we observed that only a part of the neurons are frequently employed in this task, and clear activation clusters emerge.
+This demonstrates that key-vector-based clustering can decompose modular components within the standard model.
+We then delve into expert usages across different tasks in the EMoE.
+The heatmap between tasks and experts (Fig.
+2 right) showcases that the expert utilization differs between tasks, while some crucial experts are nearly used by all 8 tasks.
+For example, expert 19, that heavily used in QQP and MNLI, is merely activated in MRPC and RTE.
+On the contrary, expert 33 is frequently activated by all 8 tasks.
+4.2 How does EMoE improve fine-tuning performance?
+Though EMoE achieves notable improvements in both ID and OOD scenarios, it is non-trivial that simply transforming the pre-trained FFNs into MoEs before fine-tuning can yield such benefits, especially when using LoRA tuning, where the gates and experts part are not updated.
+We investigate the mechanism behind the enhanced performance.
+We use BERT-Large as the backbone model.
+EMoE benefits LoRA weight learning instead of influencing inference.
+EMoE and the standard model only differ in FFNs activations.
+Such differences might (1) directly impact outputs during inference and (2) influence the parameter updating during training.
+In light of this, we propose two variants and compare them with vanilla LoRA tuning and EMoE: (a) LoRA2EMoE: Using LoRA to fine-tune the original model and split it into EMoEs during inference.
+If it surpasses the vanilla LoRA tuning, we can infer that EMoE mainly impacts model inference.
+(b) EMoE2LoRA: Using LoRA to fine-tune an EMoE model and merge experts into original FFNs during inference.
+If no changes occur, it implies that EMoE mainly ameliorates the parameter updating of the fine-tuning stage.
+Figure 3: ID and OOD accuracies compared with LoRA for validating EMoE’s training & inference effects.
+According to Fig.
+3, doing sparse activation during testing does not contribute to better generalization on average (LoRA2EMoE).
+However, when merging EMoE back into the original FFNs after fine-tuning, the performance remains significantly better than the vanilla LoRA tuning and is almost identical to EMoE (EMoE2LoRA and EMoE).
+Please refer to Appendix F.1 Tab.
+16 for full results.
+Thanks to this property, we can use EMoE during fine-tuning and then convert the models to standard ones.
+This allows the model to enjoy the benefits of EM without any alterations to its deployment, which improves EMoE’s practicability in the LLMs era.
+We further validate this on the Llama series models and the findings are consistent.
+Thus, in Tab.
+2, we report the results using EMoE to fine-tune and standard Llama architecture for benchmarking.
+And we highlight that this property can also be applied in the full fine-tuning setting, as illustrated in Appendix B.1 Tab.
+7.
+Figure 4: Sparse activated training accuracies with different expert selections.
+EMoE masks neurons with negative transfer impacts.
+The only difference between EMoE and vanilla LoRA tuning is EMoE blocks some activated neurons during training by $\operatorname{Top-k}$ expert selection.
+Based on these, we hypothesize that the effects of EMoE stem from preventing negative knowledge transfer from blocked neurons.
+Therefore, we investigate whether there are such negative transfers.
+Specifically, we study the following expert selection variants: (1) Bottom-k: select $k$ experts with the lowest scores; (2) Not-top-k: select $N-k$ experts who are not among the top-k experts.
+These variants are evaluated across six tasks from GLUE.
+The averaged outcomes are in Fig.
+4.
+Full results can be found in Appendix F.1 Tab.
+17.
+LoRA tuning results with Bottom-k and Not-top-k expert selections are worse than vanilla LoRA tuning, while Top-K outperforms it.
+We also examine the neuron activation ratio (the number of activated neurons in selected experts versus the number in the FFNs) across these variants.
+Full results can be found in Appendix F.2 Tab.
+21.
+The activation ratios for these three variants are approximately 0.43 for Top-k, 0.57 for Not-top-k, and 0.12 for Bottom-k, respectively.
+Notably, Not-top-k significantly lags behind Top-k, even though it involves and activates more neurons, suggesting the performance drop is more related to the neurons’ property.
+This further corroborates that masked neurons have negative transfer effects.
+Table 3: T5-base Multi-tasks ID and OOD accuracies across different EMoE settings.
+Experts ($N$) topk ($k$) ID avg OOD avg
+Baseline - 65.73 51.65
+8 2 67.98 52.19
+16 4 68.14 53.23
+32 8 73.29 53.20
+Figure 5: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs.
+learned) and expert types (splits of FFNs vs.
+copies of FFNs).
+The vertical axis illustrates training steps (top-down arrangement signifies begin-end); the horizontal axis represents expert selection frequency within 1K steps (deeper color implies a higher frequency).
+(a), (b) and (c) correspond to EMoE, EMoE-learn, and GMoE.
+Inspired by this, we choose a multi-task learning setting where negative transfer might be more pronounced.
+We adopt the codebase [^2] https://github.com/AkariAsai/ATTEMPT from ATTEMPT (Asai et al., 2022).
+For the ID scenario, we follow ATTEMPT and select six small tasks.
+For the OOD scenario, we train the models on two larger natural language inference (NLI) datasets and test them on four NLI datasets from different domains.
+All hyperparameters unrelated to MoEs are consistent with the baseline.
+We list the MoEs-related hyperparameters and average results in Tab.
+3.
+More details about experiment setting and results are in Appendix B.2.
+We can observe that EMoE exhibits a more substantial improvement compared to the baseline.
+In the ID setting, the highest improvement reaches 7.56, even considering the average performance across the six tasks.
+In the OOD setting, the highest average OOD result across the four datasets also improves by 1.58.
+Table 4: Results for different expert constructions (clustering-based v.s. random) and selections
+Top-k Bottom-k
+Cluster +0.92 -0.48
+Random -0.11 -0.34
+Figure 6: Results of 3 tasks for different experts splittings and expert select fractions (top-k / number of experts). ‘e=16’ and ‘e=32’ mean splitting FFNs into 16 and 32 experts, respectively.
+5 Ablation studies
+Sparsity and Modularity To provide further evidence that the EMoE’s improvements stem from leveraging modular features rather than just sparse activation or MoEs architecture, we compare the results of (1) key-vector-based clustering expert construction and (2) random construction.
+We employ the same setting of Sec.
+4.
+The relative changes in averaged outcomes compared to the vanilla fine-tuning are shown in Tab.
+4, while full results can be found in the appendix F.1, Tab.
+17.
+It’s noteworthy that while cluster top-k exhibits a significant improvement over the standard, random top-k is conversely worse than vanilla fine-tuning.
+This suggests that random construction can negatively impact gating, and only MoEs structure itself can’t bring improvement.
+Moreover, when selecting weights with negative transfer under bottom-k selections, it’s observed that cluster bottom-k also achieves lower results.
+In summary, clustering-based methods can externalize implicit modularity within the pre-trained language model.
+Within suitable frameworks like MoEs, such modularity can facilitate downstream tuning.
+MoEs Constructing Methods In EMoEs, the expert construction methods and gating methods are pivotal.
+To further understand splitting versus replicating FFNs, $\operatorname{avg-k}$ gating versus learned gating, we visualize expert selections of modularized GPT2-XL during tuning on 6 tasks with 16 experts.
+In Fig.
+5, we showcase the results for the largest dataset QNLI among them.
+Full results are available in the Appendix G.
+Our observations are: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots.
+(2) $\operatorname{avg-k}$ gating is more stable than learned gating.
+As shown in Fig.
+5 (a) and (b), the expert selection merely changes during fine-tuning.
+This might mitigate data inefficiency from gating inconsistencies across different training stages (Zuo et al., 2022a).
+(3) EMoE, with its differently initialized experts parameters, exhibits better load balancing than GMoE (expert 11 in Fig.
+5 c is way more frequently selected than other experts).
+In GMoE, all experts share identical initialization; in EMoE, the experts are derived from FFNs with EM.
+This also suggests a good initialization can facilitate MoEs learning, as indicated by Nie et al.
+(2021).
+Fig.
+5 (a) and Appendix G demonstrate that some experts in EMoE are barely selected during fine-tuning.
+This bears similarity to pruning, leading to the argument that EMoE’s improvements may stem from pruning.
+In this context, continuing our exploration from 4.2 on examining EMoE during training and inference, we supplement two ablations under the same setting: 1.
+Training-Pruning: Training EMoE (N=64, top-k=32), pruning experts with lower selection frequency, and evaluating the pruned model.
+2.
+Pruning-Training: pruning lower selection frequency experts of a new model as per step 1, then training and evaluating the pruned model.
+The experimental outcomes are detailed in Tab.
+22.
+Our primary findings are: 1.
+When sparsely selected experts are pruned post-EMoE tuning, the performance consistently outperforms the LoRA baseline (even when only 1/16 experts are used).
+This underscores that EMoE’s impact is manifested during the fine-tuning stage, not inference.
+2.
+In the optimal pruning-training setting, the performance surpasses the baseline but falls significantly short of EMoE.
+This suggests that EMoE’s improvements are not solely attributable to pruning.
+Compared with EMoE selecting experts for each token, pruning masks the same neurons for all tokens.
+Therefore, pruning can also be considered task-level EMoE, thus less effective.
+Lastly, it is important to emphasize that in multi-task learning, pruning methods struggle to ascertain the neurons requiring pruning, whereas EMoE demonstrates significant effectiveness in these scenarios.
+MoEs Configurations Beyond the settings detailed in the main results section, which are based on $N=64$ experts and top-k $\in\{16,32,48\}$, we also present specific scenarios where $N\in\{16,32\}$, and top-k varies within $\{2,4,8,16\}$ in Fig.
+6.
+Notably, in each of these settings, (1) EMoE consistently surpasses the standard model, illustrating its robustness to hyper-parameters.
+(2) On average, $\operatorname{avg-k}$ gating exhibits superior performance than learned gating.
+Though learned gating (EMoE-learn) outperforms $\operatorname{avg-k}$ gating in a few specific settings (Fig.
+6 (b) and (e)).
+This is consistent with the earlier results in Section 3.
+Regarding how many EMoE layers should be introduced, our findings align with those discovered in GMoE, indicating that only a limited number of layers can be converted into the EMoE layer.
+If excessive EMoE layers are introduced, performance deteriorates.
+Taking GPT2-XL (48 layers) as an example, when introducing EMoE every two layers in the latter half, the performance averaged across 5 GLUE tasks (79.36) matches that of the standard model (78.87).
+However, when adopting EMoE for every 2-layer for the entire model, the performance lags slightly behind that of the standard model (78.17) but surpasses the EMoE-learn (75.87).
+For additional configurations, please refer to the appendix F.2 Tab.
+19.
+6 Related Work
+Works most related to our work is introducing modularity based on off-the-shelf pre-trained models.
+For example, GMoE (Li et al., 2023) and Upcycling (Komatsuzaki et al., 2023) copy the FFNs from a trained transformer model to form the MoEs architecture.
+Their modular structure is introduced by replicating existing FFNs modules, leaving EM within pre-trained FFNs unexplored.
+MoEfication (Zhang et al., 2022b) and MoEBert (Zuo et al., 2022b) explores the EM within the model.
+They seek to improve the inference efficiency by decomposing the original FFNs layer into sparse MoEs.
+They utilize the sparse expert activation to reduce inference overhead and do not touch on how EM influences the model’s performance in downstream fine-tuning.
+Our method to split FFNs is adopted from one simple method in the ablation studies of MoEfication paper (Zhang et al., 2022a): clustering-based expert construction and $\operatorname{avg-k}$ gating.
+We empirically find that such a simple method can validate the improvements brought by EM and answer our research questions.
+We thus leave more elaborated methods for future works.
+Please refer to Appendix C for a detailed comparison of EMoE with those related works.
+7 Conclusion
+In this work, we validate that unlocking the EM in standard LMs improves downstream task ID and OOD performances.
+EMoE can bring this benefit without adding any parameters, significant training costs, or any alterations to its deployment, which improves its practicability in the LLMs era.
+One possible reason is the modular structure can alleviate negative transfer effects presented in the LMs.
+We hope our findings can deepen the understanding of neural networks’ modularity, further helping the community develop more sophisticated modular neural architectures and utilize existing LMs.
+8 Limitations
+Our primary objective was to investigate the utility of EM, and thus, we predominantly adopted the techniques from MoEfication for decomposition.
+We encourage further research to propose improved algorithms for harnessing EM.
+Our research findings have not been validated on more challenging tasks (e.g., Mathematical Reasoning (Imani et al., 2023)).
+While our analysis was primarily conducted on models with a maximum parameter count of 1.5B, we validate the scaling-up ability of EMoE to Llama-30B.
+Appendix A Additional Information
+A.0.1 Constrained clustering
+Constrained clustering (Malinen and Fränti, 2014) enhances traditional clustering methods by incorporating additional user-specified constraints.
+This approach involves balancing the k-means algorithm for clustering, making it more efficient and effective in generating meaningful clusters.
+The introduction of constraints, such as must-link or cannot-link conditions, guides the clustering algorithm to produce results that align more closely with domain-specific requirements or prior knowledge.
+This method has improved the relevancy and accuracy of clustering outcomes in various applications.
+Appendix B Additional Experiment Results
+B.1 Full Fine-tuning Performance
 Table 5: Overall OOD performances with 3 selection criteria.
@@
 Results with ViT-small (22M) backbone
-Results with ViT-base (86M) backbone
 Algorithm PACS VLCS
@@
 EMoE 87.6 79.0 73.3 45.5 71.35
+Results with ViT-base (86M) backbone
 Algorithm PACS VLCS
@@
 EMoE 89.7 79.7 77.5 48.8 73.93
-Results with ViT-base (86M) backbone
-4.1 Full Fine-tuning Performance
 We test the EMoE’s OOD performance on Domainbed.
 Domainbed provides comprehensive vision OOD evaluations (one result is aggregated with 30 experiments), and the outcomes vary marginally.
-Moreover,  Gulrajani & Lopez-Paz (2021) indicates that vanilla full fine-tuning with fair hyper-parameter search is a strong baseline compared with specifically designed methods like Invariant Risk Minimization (Arjovsky et al., 2019).
-More dataset details are in appendix A.2.1.
+Moreover,  Gulrajani and Lopez-Paz (2021) indicates that vanilla full fine-tuning with fair hyper-parameter search is a strong baseline compared with specifically designed methods like Invariant Risk Minimization (Arjovsky et al., 2019).
+More dataset details are in appendix D.0.1.
 According to Table 5: (1) Overall, EMoE outperforms ViT and GMoE (except ViT-base Train-validation setting, upper right).
@@
 Its overall performance is lower than EMoE.
-(3) In tasks where the dense model is strong (like Office), EMoE performs better than other MoEs methods.
-One possible reason is that the $\operatorname{avg-k}$ gating reduces the influence of gating weights ($g_{n}(\mathbf{x};\mathbf{G};k)$ in Equation 3), making it more like the dense model in such scenarios.
+(3) In tasks where the standard model is strong (like Office), EMoE performs better than other MoEs methods.
+One possible reason is that the $\operatorname{avg-k}$ gating reduces the influence of gating weights ($g_{n}(\mathbf{x};\mathbf{G};k)$ in Equation 3), making it more like the standard model in such scenarios.
+In vision tasks, as highlighted by (Gulrajani and Lopez-Paz, 2021), the hyper-parameter search has a profound impact on outcomes.
+Consequently, we search with a relatively large scope: for GMoE, $N$ is searched within {4, 6, 8}, and top-k within {2, 3, 4}; for EMoE, $N$ is sought within {6, 12, 24}, and top-k within {2, 4, 8}.
 Table 6: ID performance on GLUE tasks with different backbones and algorithms.
@@
 We evaluate EMoE’s ID performance on 5 GLUE tasks.
+We also include an additional baseline noisy tuning (Wu et al., 2022), which improves adaptation for free by adding uniform distribution noise proportional to the standard deviation of the pre-trained weights before tuning.
 According to Table 6: (1) On average, EMoE and EMoE-learn outperform other baselines.
@@
 (3) EMoE provides stable improvements over baselines across different settings, demonstrating its generality.
-4.2 LoRA-Tuning performance
-Table 3: ID and OOD performance of EMoE and baseline models.
-All the reported results are obtained from 3 independent experiments.
-OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization.
-The best result is highlighted in bold.
-Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
-OOD
-BERT-Large (340M Parameters)
-LoRA 89.97 63.40 72.92 90.51 93.16 92.20 87.21 85.40 84.35 4.86
-Block 89.34 62.10 71.96 90.39 93.35 92.04 88.45 86.20 84.23(-0.12) 4.95
-GMoE 89.45 63.80 72.56 90.29 93.85 92.32 87.99 85.92 84.52(+0.18) 4.04
-EMoE-block
-89.77 63.25 71.60 90.31 93.69 92.09 88.08 86.21 84.38(+0.03) 5.89
-EMoE-learn
-89.87 64.00 71.36 90.48 93.65 92.40 87.55 85.62 84.37(+0.02) 4.66
-EMoE 90.85 65.33 75.21 90.43 93.50 92.23 87.74 85.43 85.09(+0.74) 4.37
-GPT2-XL (1.5B Parameters)
-LoRA 86.83 60.88 78.70 89.07 95.18 91.84 87.41 86.93 84.61 5.61
-Block 86.59 61.18 79.78 89.08 95.45 91.88 87.71 86.95 84.83(+0.22) 5.13
-GMoE 87.02 62.81 79.78 89.21 95.41 92.18 89.10 87.17 85.34(+0.73) 4.33
-EMoE-block
-87.86 62.88 80.05 89.18 95.49 92.10 89.69 86.87 85.52(+0.91) 5.71
-EMoE-learn
-87.93 61.50 79.90 89.48 95.18 92.33 89.71 87.00 85.38(+0.77) 4.40
-EMoE 87.75 62.27 80.02 89.37 95.41 92.10 89.58 87.06 85.45(+0.84) 3.88
-With the increasing scale of pre-trained models, parameter-efficient tuning (Houlsby et al., 2019) gets popular.
-Observing weaker results with LoRA-tuned ViT-Large on Domainbed compared to ViT-Base, we omit LoRA tuning on Domainbed.
-GLUE-X (Yang et al., 2023) is a recently introduced OOD dataset for the GLUE benchmark.
-Therefore, we assess EMoE’s ID performance on GLUE and OOD performance on GLUE-X with the standard LoRA tuning.
-Notably, LoRA weights are only added in attention, with all other weights frozen.
-Therefore, the only difference between EMoE and its dense counterpart is FFNs activation.
-For the OOD metric, we follow GLUE-X (Yang et al., 2023) and employ the Friedman rank (Friedman, 1940) $\text{rank}_{f}=\frac{1}{n}\sum_{i=1}^{n}\text{rank}_{i}$.
-For each method under the same backbone, $\text{rank}_{i}$ is produced based on the best result and the average result for each dataset.
-For the 13 OOD tasks used in GLUE-X, each method generates 26 $\text{rank}_{i}$ values.
-The OOD results presented in Table 3 represent the mean of all these $\text{rank}_{i}$ values.
-The original results of each task can be found in the Appendix A.4.1.
-For algorithm configurations, since GMoE copies the FFNs of the pre-trained model to form the MoEs, it is meaningless if the MoEs parameters aren’t tuned.
-Thus, in Table 3, we conduct experiments with LoRA tuning for GMoE and also tune the transformer block where the original FFNs is replaced.
-For comparison, we also (1) LoRA-tune dense model and fine-tune the transformer block at the corresponding layer (denoted “Block”); (2) LoRA-tune EMoE with the transformed block also updated (denoted “EMoE-block”).
-According to Table 3, EMoE continues to enhance the downstream performance in LoRA tuning: (1) EMoE demonstrates significant enhancements compared to LoRA and doesn’t introduce additional trainable parameters or procedures.
-Notably, EMoE achieves comparable results with GMoE on BERT-large and outperforms it on GPT2-XL.
-(2) When the blocks are tuned, EMoE also brings improvements (EMoE-block vs.
-Block) (3) Consistent with the full-finetuning findings, EMoE exhibits higher stability than EMoE-learn and delivers superior overall results.
-(4) Besides vision, MoEs also improve OOD performance in language tasks (GMoE vs.
-Block, EMoE vs.
-LoRA).
-5 Analysis
-5.1 How does EMoE improve fine-tuning performance?
-Refer to caption     Refer to caption
-Figure 3: Left: ID and OOD results compared with LoRA for validating EMoE’s training & inference effects. Right: sparse activated training results with different expert selections.
-EMoE achieves notable improvements in both ID and OOD scenarios.
-However, it is non-trivial that simply transforming the pre-trained FFNs into MoEs before fine-tuning can yield such benefits (especially in LoRA tuning where the MoEs part is totally not updated).
-Therefore, we investigate the reasons behind the enhanced fine-tuning performance.
-To decouple the impact of updating expert parameters, we focus on LoRA tuning.
-We use BERT-Large as the backbone model.
-EMoE benefits LoRA weights learning instead of influencing inference.
-EMoE and dense model only differ in FFNs activations; such difference might (1) directly impact outputs during testing and (2) influence the parameter updating during training.
-In light of this, we propose two variants and compare them with vanilla LoRA tuning and LoRA tuning with EMoE: (a) LoRA2EMoE: splitting a LoRA-tuned dense model into its MoEs counterpart.
-If the test results differ from the original ones, we can infer that the sparse activation has an impact during testing.
-(b) EMoE2LoRA: merging the LoRA weights tuned with EMoE back into the dense.
-If no changes occur, it implies that the primary factor is the parameter updating.
-According to Figure 3 Left, doing sparse activation during testing has little impact (LoRA2EMoE).
-But when merging LoRA weights tuned with EMoE back into the dense, the performance remains significantly better than the vanilla LoRA tuning (EMoE2LoRA vs.
-LoRA) and is almost identical to EMoE.
-For full results, please refer to Appendix A.4 Table 11.
-EMoE masks neurons with negative transfer impacts.
-The only difference between EMoE and vanilla LoRA tuning is EMoE blocks some activated neurons during training by $\operatorname{Top-k}$ expert selection.
-Based on these, we hypothesize that EMoE’s effects stem from preventing negative knowledge transfer from blocked neurons.
-Therefore, we investigate whether there are such negative transfers.
-Specifically, we study the following expert selection variants: (1) Bottom-k: select $k=16$ experts who get the lowest scores; (2) Not-top-k: select experts who are not among the top-k experts.
-These variants are evaluated across 6 tasks from GLUE.
-The averaged outcomes are in Figure 3 Right.
-Full results can be found in Appendix A.4 Table 12.
-LoRA tuning results with “Bottom-k” and “Not-top-k” expert selections are worse than vanilla LoRA tuning.
-One may attribute this drop to the reduced number of employed neurons.
-To further assess this, we compare “Top-k” with “Bottom-k”, where an equal number of neurons are used.
-Notably, “Bottom-k” significantly lags behind “Top-k”.
-This further corroborates that masked neurons have negative transfer effects.
-5.2 Ablation studies
-Figure 4: Left: Activations of neurons in FFNs a pre-trained ViT (Dosovitskiy et al., 2021).
-Right: By clustering the keys in the FFNs and rearranging the action values accordingly, activation patterns of neurons emerge.
-These co-activated patterns correspond to the modular structure within the pre-trained dense model.
-With appropriate partitioning, we can transform dense networks into modular networks.
-Table 4: Clustering-based and random expert constructing results
-Top-k Bottom-K
-Cluster +0.92 -0.48
-Random -0.11 -0.34
-Expert Constructing Methods Figure 4 demonstrates that clustering can decompose modular components within the dense model.
-To provide further evidence that the EMoE’s improvements stem from leveraging modular features rather than just sparse activation, we compare the results of (1) clustering-based expert construction and (2) random construction within the same setting of Section 3.
-The relative changes in averaged outcomes compared to the dense baseline are shown in Table 4, while full results can be found in the appendix A.4, Table 12.
-It’s noteworthy that while cluster top-k exhibits a significant improvement over dense, random top-k is conversely worse than dense baseline.
-This suggests that random construction can negatively impact gating.
-Moreover, considering selecting weights with negative transfer under bottom-k selections, it’s observed that cluster bottom-k also achieves lower results.
-In summary, clustering-based methods can externalize implicit modularity within the dense model.
-Within suitable frameworks like MoEs, such modularity can facilitate downstream tuning.
-Figure 5: Results of 3 tasks for different experts splittings and expert select fractions (top-k / number of experts). ‘e=16’ and ‘e=32’ mean splitting FFNs into 16 and 32 experts, respectively.
-Expert Constructing Configurations Beyond the settings detailed in the main results section, which are based on $N=64$ experts and top-k $\in\{16,32,48\}$, we also present specific scenarios where $N\in\{16,32\}$, and top-k varies within $\{2,4,8,16\}$ in Figure 5.
-Notably, in each of these settings, (1) EMoE consistently surpasses the dense model, illustrating its insensitivity to hyper-parameters.
-(2) On average, $\operatorname{avg-k}$ gating exhibits superior performance than learned gating.
-Though learned gating (EMoE-learn) outperforms $\operatorname{avg-k}$ gating in a few specific settings (Figure 5 (b) and (e)).
-This is consistent with the earlier results in Section 4.
-Regarding how many EMoE layers should be introduced, our findings align with those discovered in GMoE, indicating that only a limited number of layers can be converted into the EMoE layer.
-If excessive EMoE layers are introduced, performance deteriorates.
-Taking GPT2-XL (48 layers) as an example, when introducing EMoE every 2 layers in the latter half, the performance averaged across 5 GLUE tasks (79.36) matches that of the dense model (78.87).
-But when adopting EMoR every 2-layer for the entire model, the performance lags slightly behind that of the dense model (78.17) but surpasses EMoE-learn (75.87).
-For additional configurations, please refer to the appendix A.4.2 Table 14.
-Figure 6: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs.
-learned) and expert types (modules from dense vs.
-repetitions of dense).
-The vertical axis illustrates training steps (top-down arrangement signifies begin-end); the horizontal axis represents expert selection frequency within 1K steps (deeper color implies a higher frequency).
-(a), (b) and (c) correspond to EMoE, EMoE-learn, and GMoE.
-Expert Selection Changes During Training To further understand $\operatorname{avg-k}$ gating and learned gating, we visualize expert selections of GPT2-XL during fine-tuning on 6 tasks with 16 experts.
-In Figure 6, we showcase the results for the largest dataset QNLI among them.
-Full results are available in the appendix A.5.
-Our observations are: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots.
-(2) $\operatorname{avg-k}$ gating is more stable than learned gating (Figure 6 a vs.
-b).
-This might mitigate data inefficiency from gating inconsistencies across different stages of training (Zuo et al., 2022).
-(3) EMoE, with its heterogeneous experts, exhibits better load balancing than GMoE (Figure 6 b vs.
-c).
-In GMoE, all experts share identical initialization, whereas in EMoE, the experts are derived from FFNs with implicit modularity.
-This also suggests a good initialization can facilitate MoEs learning similar to EvoMoE (Nie et al., 2021).
+Table 7: Instruction full-tuned Llama2-7B’s MMLU scores. Times are wall-clock computation times. The term ’times’ in the subsequent tables refers to the same concept. ’w/o’ refers to ’without’.
+Experts (N) topk (k) MMLU score times (s) FLOPS ($10^{16}$)
+w/o tuning - 46.79 - -
+full tuning - 46.5 4988 8.97
+64 16 48.08 5036 9.12
+64 32 47.44 5041 9.24
+B.2 Mult-tasks Setting
+Table 8: T5-base ID performances. All tasks are from SuperGLUE so we omit the prefix "SuperGLUE" for each tasks.
+Experts (N) topk (k) boolq cb wic wsc.fixed rte copa test avg
+Baseline 82.14 85.71 65.83 34.62 74.10 52.00 65.73
+8 2 80.67 89.29 65.52 36.54 79.86 56.00 67.98
+16 4 81.16 89.29 68.34 51.92 74.10 44.00 68.14
+32 8 80.12 78.57 70.85 63.46 82.73 64.00 73.29
+32 16 81.04 75.00 72.41 57.69 78.42 54.00 69.76
+Table 9: T5-base OOD performances. ’SG’ refers to ’SuperGLUE’.
+Experts (N) topk (k) mnli qnli wnli rte SG-rte SG-cb OOD avg
+Baseline 86.2 92.42 50.00 61.87 62.59 32.14 51.65
+8 2 86.27 92.18 50.00 64.03 62.59 32.14 52.19
+16 2 86.44 92.59 52.78 64.03 56.83 39.29 53.23
+32 8 86.56 92.49 58.33 64.03 61.87 28.57 53.20
+In our analysis 4, we have identified that the improvement brought by EMoE is likely associated with mitigating negative transfer.
+Inspired by this, we choose a multi-task learning setting where negative transfer might be more pronounced.
+We adopt the codebase [^3] https://github.com/AkariAsai/ATTEMPT from ATTEMPT (Asai et al., 2022).
+For the in-domain (ID) scenario, we follow the settings outlined in ATTEMPT and select six tasks from the Super-GLUE benchmark (Wang et al., 2019a).
+For the out-of-domain (OOD) scenario, we take two larger natural language inference (NLI) datasets MNLI and QNLI, from GLUE as our ID training data.
+We subsequently conducted direct testing on four additional NLI datasets from different domains.
+All hyperparameters unrelated to MoEs are kept consistent with the baseline, and we have listed the MoEs-related hyperparameters in the result table.
+Our observations are as follows: 1.
+EMoE exhibits a substantial improvement compared to the baseline.
+In the in-domain (ID) setting, the highest improvement reached 7.56, even considering the average performance across the six tasks.
+In the OOD setting, the highest average OOD result across the four datasets improved by 1.58.
+2.
+Across various settings of N and K, EMoE consistently outperforms the vanilla fone-tuning.
+Within the hyperparameter search space specified in our paper, EMoE consistently improves at least 2 points over the baseline in the ID setting.
+This emphasizes the effectiveness of EMoE and EMoE’s robustness to the explored hyperparameter range.
+B.3 Performance Across Different Training set Volumes
 Figure 7: Average performance of EMoE with different proportions of training data.
-Performance Across Different Training set Volumes.
 Previous research has indicated that modular architectures offer improved data efficiency (Bengio et al., 2020).
 Therefore, we conducted experiments with GPT2-XL on six tasks using varying proportions of original training data, and the results for all tasks are presented in Figure 7.
-It can be observed that EMoE consistently outperforms the dense across different data factions.
+It can be observed that EMoE consistently outperforms the standard across different data factions.
 EMoE achieves superior results even when using less than 20% of the data.
-On SST2, only using 50% data, EMoE shows comparable performance to the dense.
-More details can be found in the Appendix A.4.2 Table 15.
+On SST2, only using 50% data, EMoE shows comparable performance to the standard.
+More details can be found in the Appendix F.2 Table 20.
 This further underscores the benefits of incorporating modular structures.
-6 Conclusion, Limitations and Future Works
-Conclusion: In this work, we validate that exploiting the emergent modular structures in dense pre-trained transformers improves downstream task ID and OOD performances.
-One possible reason is the modular structure can alleviate negative transfer effects presented in the pre-trained model.
-We hope our findings could deepen the understanding of neural networks’ modularity, further helping the community develop more sophisticated modular neural architectures and utilizing existing pre-trained models.
-Limitations: We would like to point out that EMoE is not designed to outperform the SOTA tuning methods and OOD generalization algorithms.
-Instead, the goal is to demonstrate that dense pre-trained transformers can benefit from emergent modular structures.
-Our research findings have not been validated on more challenging tasks (e.g., Mathematical Reasoning (Imani et al., 2023)) or foundation models like LLaMA (Touvron et al., 2023) due to computational constraints.
-While our evaluation was primarily conducted on models with a maximum parameter count of 1.5B, it is evident that across the spectrum of parameter sizes, ranging from 22M to 1.5B, EMoE consistently yields improvements, showing the potential to scale up.
-Future works: Subsequent research could delve into better methods for modular decomposition and investigate the modularity in foundation models.
-Benefiting from EMoE’s ability to avoid negative transfer, future researchers can also exploit the emergent modularity in settings such as multi-task learning and continual learning.
-Moreover, it is observed that modularity structures emergent and remain stable after approximately one-fourth of the pre-training phase (Li et al., 2022; Zhang et al., 2023).
-Future investigations could investigate leveraging the modularity of dense models during the pre-training process.
-7 Reproducibility
-In the Experiment Configuration Section 4, we first introduce the additional configurations introduced by EMoE relative to the Dense Model, as well as the baseline settings.
-Subsequently, we provide high-level experiment settings for each experimental group (Domainbed at 4.1, full fine-tuning GLUE tasks at section 2, LoRA tuning at  3).
-In Appendix A.2, we provide detailed information on the datasets used, the codebase, and the evaluation metrics.
-We also outline more general configurations, such as learning rates and batch sizes for the respective tasks.
-Beyond the main paper, we include the original code for the experiments and log files for certain experimental results in the supplementary materials.
-Each section of code includes a README.md file that explains the experimental settings required to replicate the results.
-Appendix A Appendix
-A.1 Compare with Moefication
-Table 5: Comparison between EMoE and Moefication
-Aspect EMoE Moefication
-Motivation
-Modularity
-Sparsity
-Goal
-Transfer learning effective
-Inference efficiency
-MoEs Layers
-Typically 1 or 2 layers
-Applies changes across all layers
-Scenarios
-Down-stream tasks
-Original task
-Results
-Significant improvement without adding parameters
-Improved inference efficiency but performance drop
-Though adopting the MoEs construction method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios.
-MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons.
-In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks.
-Accordingly, We only introduce EMoEs in certain layers according to GMoE, since efficiency is not our primary concern.
+Appendix C Comparison between EMoE, MoEfication, and GMoE
+Table 10: Comparison between EMoE, MoEfication, and GMoE
+Aspect EMoE MoEfication GMoE
+Research Problem Exploit the emergent modularity during fine-tuning pre-trained transformers.
+Approximate FFNs with sparse MoE to improve inference efficiency.
+Validate the OOD improvements brought by Sparse MoE architectures.
+Methods Split FFNs Split FFNs Copy of FFNs
+Practicality No additional trainable parameters.
+Experts can be recomposed into the standard model so that models can be deployed as a standard model.
+May need re-training on the original task.
+May suffer from inference latency owing to the specific implementation of MoE architectures.
+Additional trainable parameters are introduced.
+May suffer from inference latency owing to the specific implementation of MoE architecture.
+Contribution Significant general improvement without adding parameters and not depending on the specific implementation.
+Improved inference efficiency (depending on the specific implementation of MoE), but performance drop.
+Significant OOD improvement with additional parameters and specific implementation.
+Leveraging modular designing into neural networks has various advantages (Pfeiffer et al., 2023; Chowdhery et al., 2022; Chen et al., 2023a), including superior generalization abilities (Goyal et al., 2021; Li et al., 2023).
+Most MNNs are explicitly modular.
+Among them, MoEs (Szymanski and Lemmon, 1993) is currently a standard architecture for developing MNNs (Shen et al., 2023b; Shazeer et al., 2017; Fedus et al., 2022; Zhang et al., 2022a).
+Apart from explicit MNNs,  Hod et al.
+(2021); Csordás et al.
+(2021) explore emergent modular structures in CNNs and LSTMs.
+Some recent works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated pre-trained transformers.
+Based on their observations,  Zhang et al.
+(2023) recently explore modularity in pre-trained transformer FFNs by employing handpicked semantic and knowledge-intensive tasks to probe the modular nature of pre-trained transformers.
+Having observed that FFN layers in pre-trained Transformers are sparsely activated (many neurons are unused for inputs), MoEfication splits transforms FFNs into a sparse MoE, aiming to approximate the functionality of the original FFNs to reduce the computational cost, further improving inference efficiency.
+Besides, the GMoE makes multiple replicates of the original FFN layer and introduces a learned gate to form a MoE architecture.
+They claim that such architecture could improve OOD performance from their theoretical perspective of algorithmic alignment framework.
+MoEfication and GMoE do not touch on how emergent modularity influences the training stage.
+The table below illustrates a detailed comparison of these works.
 These differences are summarized in Table 10.
-A.2 Datasets and Evaluation Metrics
-A.2.1 Domainbed
+Appendix D Datasets and Evaluation Metrics
+D.0.1 Domainbed
 Table 11: Used dataset information from Domainbed
@@
 Each dataset comprises 4 distinct domains(PACS: {art, cartoons, photos, sketches}, VLCS: {Caltech101, LabelMe, SUN09, VOC2007}, Office-Home: {art,clipart, product, real}, and Terra Incognita: {L100, L38, L48, L46}).
-Within a single training, one or two domains’ data are sequentially designated for OOD evaluation.
+One or two domains’ data are sequentially designated within a single training for OOD evaluation.
 For example, when training on PACS, {art, cartoons} could be selected as ID training data, while {photos, sketches} are designated for OOD testing.
-This configuration results in a total of $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset.
-Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited accesses to test domains and selecting based on these results.
+This configuration results in $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset.
+Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited access to test domains and selection based on these results.
 The final results are aggregated with selection criteria provided by Domainbed [^4] https://github.com/facebookresearch/DomainBed.
@@
 In our experiments, all the hyper-parameters, like training steps, learning rates, and weight decay, except those related to MoEs, strictly follow GMoE.
-A.2.2 GLUE
+D.1 GLUE
 Each task involves one to four OOD tasks from GLUE-X (Yang et al., 2023), resulting in 13 OOD tasks in total.
-To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani & Thummar, 2023) from GLUE-X for OOD testing.
+To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani and Thummar, 2023) from GLUE-X for OOD testing.
 In the full fine-tuning, to ensure convergence and reduce randomness, we train all models 10 epochs across 3 random seeds on each task.
@@
 The training batch size is 16 due to a larger model size.
-Other settings like max_lengt following the codebase from hugging face [^2] https://github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification.
+Other settings like max_lengt following the codebase from hugging face [^5] https://github.com/huggingface/transformers.
 After training on GLUE, we directly test the selected models on GLUE-X with the data from the official repo [^6] https://github.com/YangLinyi/GLUE-X.
@@
 Natural Language Inference MNLI 412,313 SICK 9,840
-A.3 Computation Cost and Memory Usage
-Theoretically, EMoE does not introduce additional parameters compared to its dense counterpart.
+Appendix E Computation Cost and Memory Usage
+E.1 Experiments with public MoEs library
+Theoretically, EMoE does not introduce additional parameters compared to the standard model.
 Although it adds computation in the gating portion within the MoEs layer, it omits a substantial amount of computation within the FFNs layer.
@@
 Since $N\ll h$, theoretically, using EMoE within a single block should accelerate the forward pass of the model.
-However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the dense model.
-Additionally, the memory usage is also slightly higher, by less than 5%, compared to the dense model (16295MB v.s.
-15865MB on CoLA, LoRA tuning GPT2-XL).
-Each individual experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
-A.4 Tabular Results
-A.4.1 Full Tables in Full Fine-tuning with Standard Deviation
-In this section, we present the mean and variance of experiments conducted with three different random seeds.
+However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the standard model.
+The memory usage is also slightly higher, by less than 5%
+Each experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
+E.2 Experiments with self-implementation
+We further find that the increasing wall time and the GPU memory usage come from the public library tutel we used to implement EMoE.
+We reimplemented our method and observed that EMoE does not require significant additional run time and memory usage.
+Specifically, we introduce an alternative implementation approach in EMoE where hidden states are used to calculate gate scores after computing the first layer.
+These scores mask the outputs of unselected experts, mimicking the effect of MoEs.
+Though this theoretically increases FLOPS compared to traditional MoEs, in practice, the speed is comparable to standard models, as demonstrated in Tab.
+72 All experiments about LLMs are conducted on eight NVIDIA 80G A100 GPU.
+Appendix F Tabular Results
+F.1 Full Tables in Full Fine-tuning with Standard Deviation
+This section presents the mean and variance of experiments conducted with three different random seeds.
 The Domainbed results are demonstrated in Table 13, full fine-tuning results are in Table 14, LoRA tuning results are in Table 15.
 Table 13: Overall out-of-domain performances with different selection criteria.
-All the reported results are obtained from three independent experiments conducted following the Domainbed codebase.
-The best result is highlighted in bold.In cases where results are the same, the best result is determined by the smallest standard deviation.
+All the reported results are obtained from three independent experiments following the Domainbed codebase.
+The best result is highlighted in bold.In cases where results are the same, the smallest standard deviation determines the best result.
 EMoE demonstrates comparable results to GMoE.
 Results with ViT-small backbone
-Results with ViT-base backbone
 Algorithm PACS VLCS OfficeHome TerraInc Avg
@@
 45.5 $\pm$0.1 71.35
+Results with ViT-base backbone
 Algorithm PACS VLCS OfficeHome TerraInc Avg
@@
 73.93
-Results with ViT-base backbone
 Table 14: Results on GLUE tasks with different backbones and algorithms.
@@
 85.38(+0.77) 3.67
-A.4.2 Full Tables and Figure Data Sources in Analysis
-In the Analysis section 5.1, for the sake of clarity, we have transformed tabular data into graphs or retained only a subset of the results.
+F.2 Full Tables and Figure Data Sources in Analysis
+In the Analysis section 4, we have transformed tabular data into graphs or retained only a subset of the results for clarity.
 The original and complete results corresponding to them are presented in this section.
-Figure 3 is summarized from Table 11 and Table 12.
-The OOD results in Table 3 are from 10.
-The ablation studies are from Table 14 and Table 15.
+Fig.
+3 is summarized from Tab.
+16 and Tab.
+17.
+The OOD results in Table 1 are from 15.
+The ablation studies are from Tab.
+19, and Tab.20.
 Table 16: ID and OOD results of BERT-L for different settings.
-”LoRA-to-EMoE” refers to converting a model tuned using standard LoRA into EMoE for testing.
-On the other hand, ”EMoE-to-LoRA” involves merging a tuned EMoE model back into a standard dense model during testing.
+"LoRA-to-EMoE" refers to converting a model tuned using standard LoRA into EMoE for testing.
+On the other hand, "EMoE-to-LoRA" involves merging a tuned EMoE model back into a standard standard model during testing.
 Algorithm CoLA Gram MRPC Twitter RTE Hans SciTail STSB Sick Avg
@@
 Table 17: ID results of BERT-L for different settings.
-”Cluster-top” refers to EMoE utilizing $\operatorname{avg-k}$ gating.
-”Cluster-not-top” represents a scenario where, during gating, the top-k experts are removed.
-Similarly, ”Cluster-bottom” involves selecting the bottom-k experts with the lowest scores during gating.
-”Random” denotes the approach of randomly selecting key values to construct experts.
-The terms ”top,” ”not-top,” and ”bottom” have the same meanings as in the cluster situations.
+"Cluster-top" refers to EMoE utilizing $\operatorname{avg-k}$ gating.
+"Cluster-not-top" represents a scenario where, during gating, the top-k experts are removed.
+Similarly, "Cluster-bottom" involves selecting the bottom-k experts with the lowest scores during gating.
+"Random" denotes the approach of randomly selecting key values to construct experts.
+The terms "top," "not-top," and "bottom" have the same meanings as in the cluster situations.
 Algorithm MRPC CoLA RTE STSB SST2 QNLI Avg
@@
 Algorithm MRPC CoLA RTE STSB Avg
-Dense 86.83
+standard 86.83
 $\pm$0.87
@@
 75.87
-Table 15: Comparison of EMoE and Dense Results with Different Training Data Fraction
+Table 20: Comparison of EMoE and Standard Results with Different Training Data Fraction
 Data CoLA MRPC RTE STSB SST2 QNLI Average
-Fraction EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense Diff
+Fraction EMoE Standard EMoE Standard EMoE Standard EMoE Standard EMoE Standard EMoE Standard Diff
 1.0 62.27 60.88 87.75 86.83 80.02 78.70 89.37 89.07 95.41 95.18 92.10 91.84 0.74
@@
 0.1 46.47 45.54 78.17 77.85 63.05 63.17 62.33 60.81 94.49 94.38 88.39 88.32 0.47
-A.5 More Visualization Results
+Table 21: Results of top and bottom selection strategies on SST2 and CoLA datasets with different activation ratios.
+The activation ratio is determined by calculating the proportion of activated neurons that belong to the selected expert among all activated neurons.
+Meanwhile, the weighted activation ratio is computed by taking the ratio of the sum of activation scores in the selected experts to the sum of the activation scores across the entire FFNs.
+Top selection Bottom selection
+Dataset Activation Ratio EMoE EMoE-learn EMoE EMoE-learn
+SST2
+Activation Ratio 32 0.6879 0.6796 0.3121 0.0552
+16 0.4317 0.4247 0.124 0.1293
+8 0.2616 0.2558 0.0528 0.3218
+Weighted Activation Ratio 32 0.731 0.7234 0.269 0.2791
+16 0.4953 0.4888 0.1041 0.109
+8 0.3304 0.3237 0.0442 0.046
+CoLA
+Activation Ratio 32 0.6815 0.6788 0.3185 0.3212
+16 0.4292 0.4239 0.1295 0.131
+8 0.264 0.2582 0.0552 0.0637
+Weighted Activation Ratio 32 0.7341 0.4958 0.2659 0.2744
+16 0.512 0.3389 0.0468 0.1109
+8 0.3589 0.7256 0.1063 0.0488
+Table 22: Comparing EMoE and pruning at different states.
+T-P means training EMoE, pruning experts with lower selection frequency, and evaluating the pruned model.
+P-T means pruning lower selection frequency experts of an untrained model, then training and evaluating the pruned model.
+CoLA SST2
+LoRA 63.40 93.16
+EMoE 65.33 93.54
+Remained Expert T-P P-T T-P P-T
+64 65.26 63.4 93.42 93.16
+32 65.33 63.52 93.54 93.34
+16 65.33 63.42 93.45 93.27
+8 64.33 63.42 93.45 93.27
+4 63.80 63.21 93.34 93.21
+Appendix G More Visualization Results
 In this part, we demonstrate more gating visualization results on SST-2, STSB, MRPC, and RTE in Figure 8.
@@
 (2) $\operatorname{avg-k}$ gating is more stable than learned gating.
-This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022).
-Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from dense vs repetitions of dense).
+This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022a).
+Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from the standard vs repetitions of the standard).
 The vertical axis illustrates training steps, with top-down arrangement signifying begin-end; the horizontal axis represents expert selection frequency within 1K steps.